### v1

# A Distributed ADMM-based Deep Learning Approach for Thermal Control in Multi-Zone Buildings

Abstract
The surge in electricity use, coupled with the dependency on intermittent renewable energy sources, poses significant hurdles to effectively managing power grids, particularly during times of peak demand. Demand Response programs and energy conservation measures are essential to operate energy grids while ensuring a responsible use of our resources This research combines distributed optimization using ADMM with Deep Learning models to plan indoor temperature setpoints effectively. A two-layer hierarchical structure is used, with a central building coordinator at the upper layer and local controllers at the thermal zone layer. The coordinator must limit the building’s maximum power by translating the building’s total power to local power targets for each zone. Local controllers can modify the temperature setpoints to meet the local power targets. The resulting control algorithm, called Distributed Planning Networks, is designed to be both adaptable and scalable to many types of buildings, tackling two of the main challenges in the development of such systems. The proposed approach is tested on an 18-zone building modeled in EnergyPlus. The algorithm successfully manages Demand Response peak events.
Note to Practitioners
This article addresses the problems faced by multizone buildings in maintaining thermal comfort while participating in Demand Response events. A centralized control approach is nonscalable, raises privacy concerns, and requires huge communication bandwidth. This work introduces a distributed algorithm that adapts the temperature setpoints of multi-zone buildings to enhance Demand Response participation and energy conservation. The proposed approach is fully data-driven and only requires historical data on weather, indoor temperature, and HVAC power consumption. The blueprint of the building and details about the HVAC system’s architecture are not needed. Energy savings will be spread across all the zones depending on the user’s choice of comfort temperature interval for each zone. The distributed optimization makes the approach scalable to large buildings with multiple zones. The algorithm is designed to work for buildings with various insulated zones such as apartment buildings and hotels, and may be extended to other types of buildings.
Index Terms:
Deep Learning, Distributed Optimization, ADMM, Demand Response, HVAC, Buildings
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Buildings account for approximately $40\%$ of the energy consumption worldwide. More than half of the building’s energy is used to maintain the occupant’s comfort, mainly through Heating, Ventilation and Air Conditioning (HVAC) systems [1]. Buildings are powerful pillars for integrating energy efficiency and Demand Response (DR) measures to reduce electricity consumption, especially during unstable power supply and peak periods. Many paths are investigated to benefit grid management by leveraging the building’s power consumption. Local Renewable Energy Sources, Energy Storage Systems, and Electric Vehicles can now be embedded in the Building Energy Management System (BEMS) to form a building-integrated microgrid [2, 3]. Efforts are also made to improve insulation materials and directly reduce the energy needs for heating and cooling purposes [4]. In addition, many grid operators are now proposing DR programs to encourage consumers to reduce their energy consumption following an incentive-based or dynamic pricing program. In both cases, they propose rewards to the end-users in exchange for reductions in their power or energy consumption, especially during periods of high solicitation of the grid. Thus, smart control of the building loads may result in substantial financial advantages for the consumers while providing support for an efficient grid operation [5, 6, 7]. Heating and cooling demands can significantly impact the total building’s power consumption profile. The thermal inertia of the building offers enhanced operational flexibility to reduce/shift a part of the load using pre-cooling or pre-heating strategies. The use of smart temperature controllers, accounting for occupancy and weather variations, results in a direct reduction of energy usage. However, it has a direct impact on the user’s comfort that must be preserved [5]. Many works have focused on the creation of advanced BEMS for monitoring and controlling a building’s energy requirements, including smart temperature controllers. Two of the main approaches used to develop BEMS are Model Predictive Control (MPC) and Reinforcement Learning (RL). For instance, [8, 9, 10] apply MPC to manage energy consumption and comfort in commercial and residential buildings. Works such as [11, 12, 13, 14, 15] use RL to create efficient and adaptive controllers for BEMS.
Optimization problems solved using MPC or RL are often formulated as centralized problems. In centralized problems, a single entity computes the solution to the problem and takes all the decisions. Centralized problem formulations for BEMS are further described in [16]. Although simple conceptually, this approach is not scalable as the system size grows [17]. Furthermore, the necessary communications between local actuators and the central controller may raise privacy concerns and failure risks. Thus, distributed architectures are increasingly considered. Distributed architectures are of two types: the hierarchical approach and the fully distributed approach (peer-to-peer). In both approaches, the complex global optimization problem is decomposed into subproblems. In hierarchical structures, the problem is solved through a coordinator who takes global decisions and communicates with Local Controllers (LCs). In fully distributed architectures, LCs directly communicate with each other to coordinate their actions. The choice of architecture is problem dependent, as each architecture has pros and cons and highly depends on the communication infrastructure [18, 11].
Distributed architectures have been increasingly used along with the Alternating Direction Method of Multipliers (ADMM) to tackle energy management challenges at the scale of the building. MPC is often considered to solve the optimization problems resulting from a decomposition, yielding a class of algorithms called Distributed Model Predictive Control (DMPC). Works using DMPC for HVAC control are further discussed in [16]. For instance, [19, 20] present case studies with linearized building dynamics, where DMPC provides equivalent performances as centralized MPC while requiring fewer computational resources. In [21], Wang et al. use a multiple-layer distributed architecture to operate HVAC systems. Their model accounts for complex coupling dynamics between the components of the HVAC system. The focus of their study is more on optimizing the operation of HVAC systems given comfort requirements. The control of the zones’ temperature setpoints to act on the power consumption has not been considered. In addition, the authors used simple linear models for the thermal dynamics of the building. In their study, Mork et al. [22] successfully implement nonlinear DMPC to control a multi-zone building. Non-linear models are seldom used in the literature but are more precise because the thermal transfer processes are non-linear [22, 23]. They use a physics-based model developed in Modelica, that accounts for the thermal and hydraulic coupling between zones. Their case studies show the efficiency of the distributed approach over centralized control. One drawback of their approach is that it requires enough knowledge about the building’s blueprint to build the model. Some data required for modelling the coupling between zones, such as door and shading positions, might also be hard to access. In addition, they apply ADMM to non-convex problems without analyzing the convergence properties of their algorithm.
On the theoretical side, ADMM is proven to converge when the underlying problem is convex [24], which is not the case in BEMS applications. To rely on the convergence guarantees of convex problems, many works linearize the dynamics of their system [19, 20, 25, 26, 27, 28]. Others rely on the empirical robustness of ADMM when applied to non-convex problems [22]. Few theoretical results exist on the convergence of ADMM in non-convex settings [29, 30]. The use of a non-convex ADMM allows to consider models more sophisticated than linear ones. In particular, models relying on Neural Networks proved to be extremely accurate to forecast the energy consumption in buildings [31, 23]. Recurrent State Space Models (RSSM) also proved to be extremely accurate to model stochastic environments and are yet to be applied to building energy forecasting. Such models, along with planning algorithms provide state-of-the-art results in many benchmark control tasks [32, 33].
In line with the International Energy Agency’s (IEA) recommendations [34], the present study focuses on the critical role of widening temperature setpoints as a way for significant energy savings and effective implementation of DR programs. While optimizing the HVAC system’s operation to efficiently track a setpoint is important, it is crucial to note that for a given temperature setpoint, there exists an inherent physical constraint on the extent of energy reduction achievable solely through the optimization of HVAC system operations. Further energy reduction may only be achieved by altering the setpoint itself. This study thus uses temperature setpoints as actuators for power consumption regulation. Within the proposed framework, users select a preferred temperature setpoint and define an associated comfort range for each zone within a building. The BEMS dynamically adjusts the setpoints within the comfort intervals to distribute energy savings across the building.
Our approach combines the recent advancement in non-convex ADMM algorithms, with state-of-the-art control methods based on Deep Learning models. First, we formulate the temperature control problem as a centralized problem with a constraint on the building’s total power consumption. Then, we convert this problem into a non-convex consensus problem that we decompose using the ADMM algorithm. We provide convergence proof of the algorithm for this non-convex setting, which is key to accurately account for the non-linear dynamics of buildings. From this formulation, we derive a surrogate problem that can be efficiently solved in practice. We test both RSSM and SSM to model the building’s zone. The models are used to plan for the temperature setpoints that best match the comfort requirements of zones while enforcing a limit on the building’s power usage. The two models are compared in terms of prediction accuracy to benchmark RSSM in a building environment, and in terms of control quality to assess the performance of our distributed control framework. The main contributions of this study are summarized as follows:
  1. 1.
    This work proposes an efficient algorithm to adjust a building’s indoor temperature based on available energy resources. The approach contributes to energy conservation while minimally affecting occupant comfort, making it a valuable asset for DR.
  2. 2.
    We combine non-convex ADMM and advanced, fully data-driven, Deep Learning models to provide a distributed control for HVAC systems. To the best of our knowledge, this is the first work combining non-linear Deep Learning models and ADMM.
  3. 3.
    While the global constraint is on the power consumption of the building, we propose to use temperature setpoints of the zones as actuators. Therefore, no assumption is made on the type of HVAC system, which makes our approach easily adaptable from one building to another.
  4. 4.
    The building is modeled using a fully data-driven approach, where no prior knowledge about the building’s blueprint is required. The RSSM allows one to plan in a stochastic environment and handle uncertainties in a robust MPC framework. The SSM can be used in computationally efficient shooting controllers.
  5. 5.
    We propose a convergence analysis for our non-convex ADMM algorithm. A case study shows the good convergence properties of our algorithm in practice.
II Problem formulation
In this study, a building composed of $N$ thermal zones is considered. Each zone has a temperature control system, i.e. a device like a PID or ON/OFF controller able to track a local temperature setpoint. Specifically, we consider a discretized setting, where occupants choose temperature setpoints $T^{sp}_{t,i}$ for each zone, that are optimal for their comfort. These setpoints may vary in time but are known beforehand. At time $t$ for the zone $i$, $u_{t,i}$ represents the power consumption of a Heating or Cooling (HC) system.
The power consumption is closely related to the temperature setpoints [35]. One can act on the setpoints to control the building’s HC power consumption below a maximum value. To this purpose, the global power constraint must be distributed to local temperature controllers acting on each zone, and translated to temperature setpoints.
The proposed control system modifies the occupants’ setpoints $T^{sp}_{t,i}$ by $\delta_{t,i}$, so that the setpoint tracked by the HC system is $T^{sp}_{t,i}+\delta_{t,i}$ at time $t$ is zone $i$. Let
$\boldsymbol{\Delta}=[\delta_{1,1},\ldots,\delta_{1,N},\ldots,\delta_{H,1},\ldots\delta_{H,N}]^{T}$
be a vector containing the setpoint changes for all the zones on a horizon of $H$ timesteps. Let $\boldsymbol{\Delta}_{t}=[\delta_{t,1},\ldots\delta_{t,N}]^{T}$ be the vector containing all setpoint changes at time $t$. Similarly, $\boldsymbol{\Delta}_{i}$ contains all the temperature setpoint changes for zone $i$ over the horizon. This indexing convention will be used for other vectors as well. In particular, the vector of power consumption:
$\textbf{u}=[u_{1,1},\ldots,u_{1,N},\ldots,u_{H,1},\ldots u_{H,N}]^{T}.$
The optimal setpoint changes $\boldsymbol{\Delta}$ are solutions to the following optimization problem:
{mini!}
[…]
where
  • $f_{i}$ represents the dynamics of zone $i$,
  • $s_{t,i}$ is the state of zone $i$. The state includes, in particular, $T^{sp}_{t,i}$, the air temperature, and the HC power consumption. The state is partially observable as it may also include the temperature of the walls or the thermal capacity of the zone.
  • $\mathcal{N}_{i}$ is the set of neighbors of zone $i$,
  • $A\in\mathbb{R}^{H\times HN}$ is a matrix that sums the power consumption of all the zones at each timestep. For instance, for two zones and a horizon of two timesteps:
    $A=\begin{bmatrix}1&1&0&0\\ 0&0&1&1\end{bmatrix}.$
    The coefficient $A_{1,1}$ selects the power consumption of zone $1$ at the first time step.
  • $\textbf{P}^{max}=[P^{max}_{1},\ldots,P^{max}_{H}]$ is a vector containing the maximum power for each timestep. The vector inequality in eq. (II) is component-wise.
  • $M_{i}$ is the maximum setpoint change defined by the occupants in zone $i$.
To solve this problem efficiently, we reformulate it as a sharing problem [29] and apply ADMM.
First, the inequality constraint (II) is relaxed using a new parameter $\textbf{P}^{tot}$ such that $\textbf{P}^{tot}\leq\textbf{P}^{max}$. This ensures the problem is feasible and removes a coupling constraint. Second, the true dynamics $f_{i}$ is approximated by a function $q_{i}$ that does not take the neighboring states as input. This assumption implies that one can reasonably predict a zone’s state without having to predict the states of neighboring zones. Note that the initial state of a zone $i$, $s_{0,i}$ may still be augmented to include information about the neighbors at $t=0$. In many buildings such as hotels, offices, or residential buildings, the temperature ranges in the zones are similar, thus limiting the heat exchanges. We will assess this assumption experimentally in section V to show that it is reasonable if the zones are well insulated. Such an assumption is also discussed in [22, 36]. The optimization problem becomes:
{mini!}
[…]
The HC power consumption is now driven towards the parameters $\textbf{P}^{tot}=[P_{1}^{tot},\ldots,P_{H}^{tot}]$. The choice of $\textbf{P}^{tot}$ is important and will be further discussed in section V.
Let us define a function $\phi_{i}$ that maps the state of a zone $i$ to its power usage:
$u_{t,i}=\phi_{i}(s_{t,i}).$
Given the setpoints and the first state, future states are computed by unrolling the dynamics:
$s_{1,i}$ $=q_{i}(s_{0,i},\delta_{0,i})$
$s_{2,i}$ $=q_{i}(s_{1,i},\delta_{1,i})$
$=q_{i}(q_{i}(s_{0,i},\delta_{0,i}),\delta_{1,i})$
$\ldots$
$s_{H,i}$ $=q_{i}(s_{H-1,i},\delta_{H,i})$
$=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{H,i})$
To ease the notations, we define the following functions, for all $i$ and $t$:
$Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{t,i})$ (1)
$g(\boldsymbol{\Delta}_{i})=\left\lVert\boldsymbol{\Delta}_{i}\right\rVert^{2}_{2}$ (2)
$\ell(\boldsymbol{\Delta})=\sum_{t=1}^{H}\left(\sum_{i=1}^{N}\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})-\textbf{P}^{tot}_{t}\right)^{2}$ (3)
The index to the function $Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})$ indicates that it represents the state at time $t$ of zone $i$ computed from $s_{0,i}$. Only the first $t-1$ elements of $\boldsymbol{\Delta}_{i}$ are needed to compute $s_{t,i}$, but the entire vector is passed for notational simplicity. The power usage becomes:
$u_{t,i}=\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i}))$ (4)
[…]
This problem is a sharing problem as described in [29]. Each zone has a local objective represented by $g$ and a global objective represented by $\ell$. The total power available is a common resource that must be shared across all zones to satisfy the local comfort requirements.
[…]
The matrices $B_{i}\in\mathbb{R}^{NH\times H}$ for $i=1,\ldots,N$ are such that $\boldsymbol{\Delta}=\sum_{i=1}^{N}B_{i}\boldsymbol{\Delta}_{i}$. The augmented Lagrangian, with Lagrange multipliers $\boldsymbol{\lambda}_{1},\ldots,\boldsymbol{\lambda}_{N}\in\mathbb{R}^{H}$ associated with the equality constraints, is:
$\mathcal{L}_{\rho}(\boldsymbol{\Delta},\bar{\boldsymbol{\Delta}},\boldsymbol{\lambda}_{1},\ldots,\boldsymbol{\lambda}_{N})=\sum_{i=1}^{N}g(\boldsymbol{\Delta}_{i})+\ell(\bar{\boldsymbol{\Delta}})$
$+\sum_{i=1}^{N}\left(\boldsymbol{\lambda}_{i}^{T}(\bar{\boldsymbol{\Delta}_{i}}-\boldsymbol{\Delta}_{i})+\rho\|\bar{\boldsymbol{\Delta}_{i}}-\boldsymbol{\Delta}_{i}\|^{2}_{2}\right).$ (5)
Algorithm 1 :non-convex consensus ADMM
1:  Initialize $\text{Initialize }\{\boldsymbol{\Delta}_{i}\}_{i=1,\ldots,N},\{\bar{\boldsymbol{\Delta}}_{i}\}_{i=1,\ldots,N},\{\boldsymbol{\lambda}_{i}\}_{i=1,\ldots,N}$
2:  $k=0$
3:  while has not converge do
4:     for $i=1,\ldots,N$ do
5:        $\Delta_{i}^{k+1}\leftarrow\operatorname*{arg\,min}_{\boldsymbol{\Delta}_{i}}g(\boldsymbol{\Delta}_{i})+\boldsymbol{\lambda}_{i}^{k,T}(\bar{\boldsymbol{\Delta}}_{i}^{k}-\boldsymbol{\Delta}_{i})\newline \hskip 71.13188pt+\frac{\rho}{2}||\bar{\boldsymbol{\Delta}}_{i}^{k}-\boldsymbol{\Delta}_{i}||^{2}_{2}$
6:     end for
7:     $\bar{\boldsymbol{\Delta}}^{k+1}\leftarrow\operatorname*{arg\,min}_{\bar{\boldsymbol{\Delta}}}\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{\Delta}}_{i}\right)\newline \hskip 14.22636pt+\sum_{i=1}^{N}(\boldsymbol{\lambda}_{i}^{k,T}(\bar{\boldsymbol{\Delta}}_{i}-\boldsymbol{\Delta}_{i}^{k+1})+\frac{\rho}{2}||\bar{\boldsymbol{\Delta}}_{i}-\boldsymbol{\Delta}_{i}^{k+1}||^{2}_{2}$ )
8:     for $i=1,\ldots,N$ do
9:        $\boldsymbol{\lambda}_{i}^{k+1}\leftarrow\boldsymbol{\lambda}_{i}^{k}+\rho(\bar{\boldsymbol{\Delta}_{i}}^{k+1}-\boldsymbol{\Delta}_{i}^{k+1})$
10:     end for
11:  end while
The sharing problem can be solved using Alg. 1. We distinguish two main sub-problems in this algorithm. The first one is the local controller problem (line 5) and the second one is the coordinator problem (line 7).
As stated in the following theorem, under some Lipschitz regularity assumptions and the right choice of $\rho$, this algorithm is guaranteed to converge although the problem is not convex. We will assess the quality of the solutions experimentally in section V.
Theorem II.1.
The Alg. 1 converges to the set of stationary solutions of problem II :
$\lim_{k\xrightarrow{}\infty}dist(\{\boldsymbol{\Delta}_{i}^{k}\},\{\bar{\boldsymbol{\Delta}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\};Z^{*})=0,$
where $Z^{*}$ is the set of primal-dual stationary solutions of the problem.
Evaluating the models of each dynamics $q_{i}$ is computationally expensive. To leverage the benefits of the decomposition, the models’ evaluation should be placed in the for loop and computed in parallel. In Alg. 1 the models’ evaluation is in the $\ell$ function, in the coordinator problem.
We can explicit the relationship between the setpoints and the related power consumption (4) as a constraint in the LC problem.
{mini!}
[…]
and consider the surrogate coordinator problem:
{mini!}
[3] ¯u~ℓ(¯u) + ∑_i=1^N λ_i^T (¯u_i - u_i) + ρ—— ¯u_i - u_i ——_2^2
where ~$\tilde{\ell}(\bar{\textbf{u}})=\left\lVert A\bar{\textbf{u}}-\bar{P}^{tot}\right\rVert$. This formulation may be derived from the same decomposition steps as the ones above, but using the following equation rather than eq. (4):
$\delta_{t,i}=\psi(Q_{t,i}(s_{0,i},u_{i})),$ (6)
Eq. (6) bounds the state and power usage to a setpoint. The problem is that $\psi$ is not an injective application. For instance, if the temperature in a zone is $20^{\circ}C$, temperature setpoints below $20^{\circ}C$ will all yield a heating power of $0W$. Such an application is impossible to learn accurately in practice. That said, depending on the type of setpoint tracking system, how the power is measured and the operation regime, one may find conditions for $\psi$ to satisfy the requirements of Theorem 1. For instance, a proportional controller that operates with a zone’s temperature around the setpoint or an ON/OFF controller used with Pulse Width Modulation would give good properties to $\psi$.
Instead of enumerating such conditions, that are hard to verify in practice, we choose to provide a theoretical analysis for the problem formulation (4). This formulation is more natural and indicates the form of the ADMM decomposition for the surrogate problem. As shown in Section V, the surrogate formulation yields efficient computation and shows excellent convergence properties experimentally. In addition, it provides a good intuition about the underlying mechanism: the coordinator (i.e. problem (II.1)) is choosing the amount of power for each room and each timestep, to ensure that the power constraint is always satisfied. The LCs (i.e. problems (II.1)) must then find the temperature setpoints to track this power target while preserving the users’ comfort. Even if multiple setpoints may satisfy the constraints II.1 and II.1, the lowest setpoint changes will be chosen because of the term $g(\Delta_{i})$ is the objective function.
The control architecture is summarized in Fig. 1.
Figure 1: Two layer hierarchical control architecture. The LCs observe the state of their zone and plan for the best setpoints while communicating with the Coordinator to enforce a maximum constraint on power usage. Setpoints are sent to the existing setpoint trackers that act on each zone.
III Local Controllers and Zone Modelling
Figure 2: Distributed Planning Network - For each iteration of the ADMM algorithm, the coordinator sends power constraints to the LCs. Using the last observation, the LCs plan for the best trajectory of actions, either by shooting or random search, and send the predicted powers to the coordinator. Once the ADMM has converged, only the first action of the horizon is applied to the zones.
To successfully plan for the best setpoint changes without implementing any actions, LCs must be able to predict the evolution of the system. In this section, we present two methods for the modeling of a thermal zone: Recurrent State Space Models (RSSM) and State Space Models (SSM).
III-A Reccurent State Space Model
The RSSM is a stochastic model that allows forecasting of the power usage of each zone under uncertainties brought by many unobservable disturbances. The use of RSSM for planning is further described in [32]. In this section, we focus on some key elements for completeness.
The RSSM is a black-box model, learned from previously collected data. It is composed of three main parts:
  1. 1.
    A representation model (encoder), which maps the observations $o_{t}$ to the latent state $s_{t}$ of the system. The use of an encoder is important because the system is partially observed. One can include lags of power and temperature in the observation to build a better representation of the state and cope with partial observability.
  2. 2.
    A transition model, that predicts the evolution of the latent space. Note that this transition model is composed of two parts: (1) a deterministic state model which predicts the evolution of the deterministic part of the state $h_{t}=f(h_{t-1},s_{t-1},\delta_{t-1})$ and (2) a stochastic state model which samples the next latent state according to $h_{t}$ : $s_{t}\sim p(s_{t}|h_{t})$. The latter part is important in our case as it accounts for the perturbations due to the occupancy and neighboring zones.
  3. 3.
    An observation model (decoder), that outputs the observation corresponding to a given latent state.
All these components are represented by neural networks. We use the same set of parameters $\theta$ to learn the observation model, the transition model, and the representation model. The objective is to maximize the variational bound as proposed in [32]. The transition model allows to predict one timestep ahead in the latent space. Multiple steps prediction is achieved by applying the model recursively. This is convenient as the user may want to change the prediction horizon without having to retrain the model. The latent space embeds the inner state of the room, mainly composed of the power usage and the room’s temperature. This state is influenced by exogenous perturbations such as weather and the temperature setpoint. To not overload the transition model, we do not include these disturbances in the latent space as it is usually done with this type of model. Instead, we feed the transition model with the previous latent space as well as forecasts of the disturbances, as pictured in Fig. 3. The advantages of this approach are that a single weather prediction system is required for the entire building. Since the RSSM of each zone does not have to predict the complex evolution of the weather on top of the evolution of the room’s state, it can provide better predictions using a smaller network model. In addition, some disturbances such as temperature setpoint schedules may be known beforehand and can be passed exactly to the transition model at each prediction step.
Figure 3: Multi-step ahead power predictions with weather forecasts
III-B State Space Models
The State Space Model (SSM) has the same structure as the RSSM, but it does not include the stochastic part. The transition model is a GRU cell that takes as input the disturbances (i.e. the weather and the setpoints). The latent state $s_{t}$ is carried by the hidden state of the GRU cell. The model is trained to minimize the mean squared error between the predicted and actual powers.
IV Distributed Planning Networks
In this section, we present how the models of each LC are combined with the coordinator using ADMM. The proposed algorithm is called Distributed Planning Networks as it embeds the idea of the Planning Networks in a distributed architecture. Fig. 2 and Alg. 3 summarize the approach.
The coordinator problem is a convex penalized least square problem. It can be solved easily with existing solvers. The challenge is more on the side of the LC problem (II.1).
To efficiently find a solution to the non-convex LC problem, we propose two approaches. The first one is a random search for the best temperature setpoints. The second one is a shooting method, where we differentiate through the prediction model and use the gradient to guide the search. Once the best setpoint changes have been found for the upcoming horizon, only the setpoint changes of the first timestep are implemented. The process is then repeated to allow for re-planning and correction of the control.
IV-A Planning by Random Search with RSSM
The random search is conducted as follows:
  1. 1.
    Generate random sequences of setpoint changes $\boldsymbol{\Delta}_{i}$.
  2. 2.
    Sample the trajectories corresponding to each $\boldsymbol{\Delta}_{i}$ using the RSSM. For each $\boldsymbol{\Delta}_{i}$, the RSSM outputs a distribution for the trajectory of states. For one choice of $\boldsymbol{\Delta}_{i}$, multiple states trajectories may be sampled from the distribution to account for uncertainties.
  3. 3.
    Select $\boldsymbol{\Delta}_{i}$ providing the lowest score as defined by (II.1).
While using ADMM, the steps presented above need to be repeated at each iteration. The second step is computationally expensive as it requires many forward passes in the transition and observation networks. However, only the third step depends on the ADMM iteration. To limit the computational time, the trajectories (i.e. the two first steps) are pre-computed and stored before the start of ADMM. Note that these two steps can be performed in parallel for each zone. At each ADMM iteration, the score of each trajectory is then computed from the pre-computed trajectory with respect to the current power targets $\bar{\textbf{u}}_{i}$.
IV-B Planing by Shooting with SSM
The SSM is fully differentiable. One can formulate the problem (II.1) as an Initial Value Problem and use single shooting to minimize the running cost given by the objective function. The equality constraints representing the dynamics are implicitly satisfied by the prediction model. To handle the inequality constraint (II.1), we use projected gradient descent to keep the setpoint changes within the desired interval. This approach assumes continuous values for the setpoints. In practice, we round the values at the resolution of the setpoint tracker of the zone afterwards. A stopping criterion on the number of decimals changing in $\Delta_{i}$ at each iteration is used to make the approach computationally efficient.
V Results
To test the proposed approach, we use the low-rise apartment building, one of the Department Of Energy (DOE) residential reference buildings. The building is modeled in EnergyPlus, a software that provides thermal simulations for buildings with state-of-the-art accuracy. The building is composed of three floors of 6 apartments each. Apartments are equipped with independent HVAC systems, temperature setpoints schedules, occupancy schedules, and appliance usage schedules. Simulations are run using Typical Meteorological Year (TMY) files coming from the DOE climate zone 7A, representing the weather in the North-East of the United States of America and South-East of Canada. In each apartment, the user setpoint $T_{i}^{sp}$ is kept constant, but may vary between $19^{\circ}C$ and $22^{\circ}C$ from one apartment to the other. For the purpose of this study, we are considering heating periods, and hence we modify the default HVAC systems to electric heating-only systems with an efficiency of one. The EnergyPlus timestep was set to 15 minutes. The prediction horizon is set to one or two hours and the ADMM timestep is 30 minutes, meaning we can re-plan to change the setpoints every 30 minutes.
V-A Models accuracy
zone RSSM SSM zone RSSM SSM
1h ahead forecast 1 185 78 10 125 89
2 129 79 11 158 140
3 181 88 12 127 71
4 136 67 13 177 92
5 122 55 14 134 94
6 130 68 15 164 105
7 137 94 16 152 69
8 144 116 17 128 57
9 159 107 18 134 84
zone RSSM SSM zone RSSM SSM
2h ahead forecast 1 184 80 10 98 82
2 120 71 11 144 133
3 182 88 12 109 72
4 134 63 13 173 103
5 104 51 14 112 96
6 127 66 15 159 110
7 127 105 16 122 71
8 131 121 17 105 79
9 152 108 18 109 102
TABLE I: Models scores (MAE, Watt) for one hour and two hours ahead predictions during a three weeks test period. For the entire building’s power forecast, the RSSM has a MAPE of $5.81\%$ for one hour ahead and $6.20\%$ for two hours ahead. The SSM has a MAPE of $3.72\%$ and $4.58\%$ respectively.
[…]
First, we test the accuracy of the prediction models. Both RSSM and SSM are trained offline for each zone of the building. The training data are generated by running simulations with random temperature setpoints changed every period of random duration, ranging from 1h to 24h. The observation of a zone contains the zone’s current temperature as well as the power consumption of each timestep for the last hour. The output of the representation model is the power consumption of the next timestep.
At each timestep, the zone’s observation is completed with disturbances inputs. Disturbances include the temperature setpoint and the temperature setpoint changes that are known beforehand. As one may observe in Fig. 4, the changes in the temperature setpoints cause steep variations in the power. To accurately predict these changes, a key feature added is the difference between setpoint changes from one timestep to the next i.e. $\delta_{t,i}-\delta_{t-1,i}$. The disturbances also include the weather, composed of the outdoor temperature and solar radiation. For simplicity, we did not build a weather forecast model. Instead, we used perfect weather information and added random noise to account for prediction errors.
Table I shows the Mean Absolute Error (MAE) of each zone for one-hour and two-hours ahead forecasts. The SSM performs consistently better than the RSSM, but both methods give accurate predictions. For the entire building’s power forecast, the Mean Average Percentage Errors (MAPE) of the SSM are $3.72\%$ and $4.58\%$ for one-hour and two-hours ahead predictions. The MAPE of the RSSM are $5.81\%$ and $6.20\%$ respectively. These predictions were made without observing the states of the neighboring zones, as stated in the assumption of the model used in (II).
V-B Distributed Planning Networks and Power Control
To test the Distributed Planning Networks algorithm, we propose to implement it for peak load reduction. We consider a Demand Response (DR) program where the utility sends an early notification to the building’s operator to reduce the power demand during peak demand events. We assume that the events occur between $5AM$ and $9AM$ every morning over a period of three weeks. We also assume that this reduction implies a limit on the heating demand that should not exceed $\textbf{P}^{max}_{t}=40kW$ for all timesteps $t$ during DR events. For each zone, we set $-2^{\circ}C\leq\boldsymbol{\Delta}_{t}\leq 0^{\circ}C$, with an action step of $0.25^{\circ}C$.
Algorithm 2 Constraint conversion from $\textbf{P}^{max}$ to $\textbf{P}^{tot}$
1:  Inputs : current state $\text{Inputs : current state }s_{t}\text{ and }\textbf{P}^{max}$
2:  Prediction models $\text{Prediction models }\phi_{1},\ldots,\phi_{N}$
3:  Parameters : $\text{Parameters : }\nu,\boldsymbol{\Delta}_{min},\boldsymbol{\Delta}_{max},H$
4:  Initialize $\textbf{P}^{tot}\in\mathbb{R}^{H}$
5:  for i=1, …, N do
6:     $P_{normal,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{max}))$
7:     $P_{min,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{min}))$
8:  end for
9:  $\textbf{P}_{normal}=\sum_{i=1}^{N}P_{normal,i},\text{ }\textbf{P}_{min}=\sum_{i=1}^{N}P_{min,i}$
10:  if $\textbf{P}_{normal}\leq(1-\nu)\textbf{P}^{max}$ then
11:     $\delta_{t,i}=0\text{ }\forall t,i$
12:     DO ADMM = FALSE
13:  else if $\textbf{P}_{min}>(1-\nu)\textbf{P}^{max}$ then
14:     $\delta_{t,i}=-2\text{ }\forall t,i$
15:     DO ADMM = FALSE
16:  else
17:     $\textbf{P}^{tot}=clip(\textbf{P}_{normal},(1-\nu)\textbf{P}^{max})$
18:     DO ADMM = TRUE
19:  end if
As mentioned in section II, translating the constraint’s parameter $\textbf{P}^{max}$ to $\textbf{P}^{tot}$ (defined in eq. (II) ) is important as the building power consumption will be driven towards $\textbf{P}^{tot}$. To find the right $\textbf{P}^{tot}$, we propose to leverage the prediction model of each zone. We forecast the total power for setpoint changes of $0^{\circ}C$ and $-2^{\circ}C$ to find lower and upper bounds on the power. If the lower bound is below $\textbf{P}^{max}$, the optimal solution of the problem II is $\boldsymbol{\Delta}=0$ and no further computation is required. If the upper bound is above $\textbf{P}^{max}$, the problem (II) is unfeasible. In the latter case, depending on the energy pricing, one could either reduce as much as possible the temperatures to lower the consumption, or take no action not to alter the comfort since the DR requirement cannot be met. If $\textbf{P}^{max}$ falls between the upper and lower bounds, we proceed with the Distributed Planning Networks algorithm. We then set $\textbf{P}^{tot}=\textbf{P}^{max}$. Note that in practice, we use a slack parameter $\nu$ to mitigate the effects of prediction errors, and set $\textbf{P}^{tot}=(1-\nu)\textbf{P}^{max}$. This procedure is detailed in Alg. 2.
V-B1 Planning by shooting with SSM
During the three week test period, the power consumption was successfully reduced below $\textbf{P}^{max}$ during each DR event. Fig. 5 shows the power consumption during three days of the test period.
Fig. 6 shows the effect of the slack parameter $\nu$ on the heating power consumption. The more $\nu$ increases, the more conservative the control is but the higher the chance to satisfy the power constraint despite prediction errors. In our experiment, the power constraint was satisfied for the three choices $\nu=8\%$, $\nu=4\%$, and $\nu=0\%$. However, the maximum power monitored during DR event increased as the $\nu$ decreased. With $\nu=0\%$ the power reached $39.9kW$, compare to a maximum of around $38.5kW$ for $\nu=8\%$, making the power constraint really tight. The right choice of $\nu$ is therefore a trade-off between comfort and power usage.
In Fig. 5, we observe rebound peaks caused by the increase of temperature setpoints after the DR events. These peaks may be managed in several ways. For instance, as shown in Fig. 7, one may extend the duration of the DR event to reduce the magnitude of the rebound peaks. Our algorithm may also be used for peak shaving by maintaining the power constraint throughout the day, as shown by the green curve in Fig. 7.
Table II shows the distribution of the temperature setpoints during active DR events. We observe that the effort is evenly spread among the zones. No zone had to take an extreme $-2^{\circ}C$ action and most of the actions are between $-0.5^{\circ}C$ and $\-1^{\circ}C$. Note that we did not enforce any type of fairness in the algorithm. We initialize the ADMM iterations with a setpoint change of $-1^{\circ}C$ and the same power target for every zone. We attribute this equal distribution of actions to the initialization. Differences arise because each zone has a different heating power requirement. Power consumption is influenced by many factors such as the orientation of the building, the position of the zone within the building, the presence of windows, and the occupancy.
V-B2 Planing by random search with RSSM
SSM zone mean $(^{\circ}C)$ None Low Medium High zone mean $(^{\circ}C)$ None Low Medium High
1 -0.72 0 34.21 65.79 0 10 -0.61 0 52.63 47.37 0
2 -0.56 0 65.79 34.21 0 11 -0.74 0 44.74 47.36 7.90
3 -0.80 0 18.42 81.58 0 12 -0.78 0 26.31 71.05 2.64
4 -0.73 0 34.21 65.79 0 13 -0.84 0 15.79 78.95 5.26
5 -0.66 0 42.10 57.90 0 14 -0.61 0 55.26 44.74 0
6 -0.78 0 21.05 78.95 0 15 -0.78 0 26.32 71.05 2.63
7 -0.70 0 39.48 57.89 2.63 16 -0.82 0 18.42 78.95 2.63
8 -0.66 0 50.0 44.73 5.27 17 -0.79 0 21.06 76.31 2.63
9 -0.74 0 39.47 55.26 5.27 18 -0.73 0 26.32 71.05 2.63
Model RSSM zone mean $(^{\circ}C)$ None Low Medium High zone mean $(^{\circ}C)$ None Low Medium High
1 -1.14 8.12 13.51 24.32 54.05 10 -1.29 5.41 13.51 13.51 67.57
2 -1.27 2.72 13.51 24.32 59.45 11 -1.32 2.70 8.11 18.92 70.27
3 -1.24 2.71 10.81 29.72 56.76 12 -0.9 13.52 35.14 8.10 43.24
4 -1.14 8.12 8.11 24.32 59.45 13 -1.17 8.12 21.62 13.51 56.75
5 -1.35 0 18.92 16.22 64.86 14 -1.32 2.70 13.52 13.51 70.27
6 -1.41 0 8.11 13.51 78.38 15 -0.78 0 51.35 16.22 32.43
7 -1.23 0 16.21 24.32 59.46 16 -1.16 5.41 18.92 16.22 59.45
8 -1.36 0 16.22 10.81 72.97 17 -1.17 0 24.33 21.62 54.05
9 -1.41 0 8.10 16.22 75.68 18 -1.15 0 24.32 24.32 51.36
TABLE II: Setpoint changes during active DR events (i.e. DR events where actions were required). The values correspond to the percentage of time spent with the following setpoint change : None := $\delta=0$, Low := $0.50<\delta\leq-$ 0.5, Medium := $-0.5<\delta\leq-1.0$, High := $-1.0<\delta\leq-1.75$. The results for the SSM are obtained with $\nu=4\%$.
A random search is used along with the RSSM to plan for the best setpoints. We split the computational budget between exploring the action space (i.e. sampling trajectories for different $\boldsymbol{\Delta}_{i}$) and exploring the possible trajectories for each choice of $\boldsymbol{\Delta}_{i}$. At the beginning of each timestep, we randomly sample 100 different choices for $\boldsymbol{\Delta}_{i}$, and for each $\boldsymbol{\Delta}_{i}$, we sample 25 different trajectories. We then select the trajectory out of the 25 that yields the highest power consumption, to optimize based on the worst-case scenario. Fig. 8 shows an example of different power consumptions sampled from a given state for a choice of $\boldsymbol{\Delta}_{i}$. Note that with this approach, we fix the slack parameter $\nu$ to zero because the optimization is already made based on a worst-case scenario. The experiment is repeated for 5 different seeds. The results are consistent across all the seeds. For simplicity, only one outcome is presented.
Using this type of controller, the power is successfully controlled below $P^{max}$. The maximum power monitored during a DR event is $39.54\pm 0.56kW$  [^1] Mean and standard deviation computed on the different seeds., resulting in a tight power constraint. To build a better estimate of the worst-case scenario, one can increase the number of trajectories sampled for each $\Delta_{i}$. Alternatively, one may choose a strictly positive slack parameter to mitigate the impact of prediction errors on the satisfaction of the maximum power constraint. Fig. 5 shows a three days example of power consumption. Note that the rebound peaks may also be managed by increasing the duration of DR events.
As shown in Table II, the effort is spread across all the zones. As for the shooting method, the algorithm is initialized evenly by fixing the same power targets for every zone. As a result of the optimization based on the worst-case scenario, the temperature setpoint reductions are higher than with the shooting method. This results in a larger power reduction during DR events, as seen in Fig. 5.
One of the main limitations of this approach is the exploration of the action space. For a one-hour ahead planning, the action space is small and may be well explored by random search. However, to plan on a longer horizon, it is possible to reallocate the computational budgets to sampling more $\boldsymbol{\Delta}_{i}$ rather than different trajectories for a given $\boldsymbol{\Delta}_{i}$. The limiting factor is time, as the computation must be made within the $15$ minutes timestep. In our experiments, on a single CPU (3.5GHz), the sampling of all the $2500$ trajectories took around $5$ minutes. This computation is done in parallel for all the zones. Once the trajectories have been sampled, a hundred ADMM iterations can be performed within a minute.
Figure 5: Comparison between the random search and the shooting control. Demand Response events occur between $5$AM and $9$AM every day. The slack is set to $\nu=4\%$ for the shooting control.
Figure 6: Influence of the slack parameter $\nu$ on the power consumption for the shooting control.
Figure 7: Influence of the duration of the DR event on the rebound peak. The black line represents $\textbf{P}^{max}$. Our algorithm is flexible and can also be used for peak shaving.
Figure 8: Different power trajectories for a zone, sampled from a given state $\boldsymbol{\Delta}_{i}$ using a RSSM
V-B3 Convergence Analysis
Figure 9: Mean and standard deviation of the residuals for ADMM and random search with a RSSM or shooting with a SSM. The statistics are taken over the first $30$ iterations of all the ADMM usage over the test period.
Fig. 9 shows the mean and standard deviation of the residuals, taken over the first $30$ iterations of each ADMM usage. In both cases, the residuals are decreasing and the algorithm shows excellent convergence properties.
The algorithm converges rapidly in the first few iterations. In practice, we found out that performing just 20 iterations of ADMM is a good choice to save computational time. It makes little difference when using RSSM because trajectories are sampled beforehand. However, with the shooting method, the computation time increases linearly with the number of ADMM iterations. It takes one minute to perform ten iterations and ten minutes to perform a hundred iterations on a machine with 8 CPUs (3.5GHz).
VI Discussion and further work
The results presented above demonstrate the effectiveness of the algorithm to distribute a global coupled power constraint to local heating systems in a multi-zone building.
The experiments have revealed that a low prediction error is key to the success of the method. Prediction errors are not an issue if one aims at maintaining the power around a target value. However, most of the constraints implying a maximum power are hard constraints and an inaccurate forecast will cause the constraint to be violated. The maximum prediction error could easily exceed the mean error of $3.72\%$ or $5.81\%$ reached by our models, especially at the end of peak periods where resetting the setpoint changes to zero causes sharp variations of the power usage. To improve the accuracy of the models, one may include the state of the neighboring zones in the observations of a zone. Another possibility is to retrain the models on data collected during the control period. Such data are more representative of encountered states than the data coming from random control of the temperature setpoints used to first train the models.
An interesting extension of this work would be to measure thermal comfort based on predictions of the temperature of each zone instead of on the temperature setpoints. The local objectives $g$, defined in eq. (4), would become
$g(\boldsymbol{\Delta}_{i})=\left\lVert T_{t,i}-T^{sp}_{t,i}\right\rVert_{2}^{2},$
where $T_{t,i}$ is the temperature in zone $i$ at time $t$. This complicates the theoretical analysis because such a $g$ function is no longer convex. In addition, some preliminary experiments suggest that it is hard to accurately predict the temperature with fully data-driven models. If the temperature setpoint is passed as an input, the models often learn to predict the setpoint as the temperature forecast. The use of physics-informed models may be necessary to ensure accurate temperature predictions [23, 37].
We trained all the prediction models from scratch. However, there are many similarities from one zone to the other. An interesting improvement for this work will be to leverage the training of a model in one zone to accelerate the training of other models. Existing techniques of Meta-Learning could be used for this purpose. This will greatly improve the generalization of the approach as models from some buildings could be used for new buildings.
Finally, tests are all performed using a virtual environment. Even though EnergyPlus provides state-of-the-art accuracy for building simulation, future works should include tests on real buildings to further validate the method.
VII Conclusion
In this work, we present a distributed algorithm that combines ADMM and Deep Learning models to control the temperature setpoints and act on the power consumption of HVAC systems in buildings. Based on a global coupled power constraint, the algorithm distributes the power across the Local Controller systems. This approach does not assume any particular type of local heating or cooling system and learns its model entirely from data. We provided a theoretical framework and derived a practical algorithm that we implemented to test the efficiency of the approach. Tests are performed on a large residential building composed of 18 zones, modeled using EnergyPlus. By combining the zones’ prediction models, we reached a mean error of $3.72\%$ with SSM and $5.72\%$ with RSSM on the prediction of the total heating power consumption. The Distributed Planning Networks algorithm is tested on a power control task during DR events. The power is successfully maintained below a given maximum power limit and the effort is shared across all zones.
Appendix A Appendix
A-A Distributed Planning Networks
Algorithm 3 Distributed Planning Networks
1:  $N\text{number of zones, }H\text{ prediction horizon}$
2:  $\{LC_{i}\}_{i=1,\ldots,N}\text{ Local Controllers}$
3:  $\{ST_{i}\}_{i=1,\ldots,N}\text{ Setpoint trackers}$
4:  $Agg\text{ coordinator}$
5:  $Env\text{ Building Environment}$
6:  $\{\mathcal{D}\}_{i=1,\ldots,N}\text{ collect data from }Env$
7:  Train $\text{Train }LC_{i=1,\ldots,N}\text{ on }\mathcal{D}_{i=1,\ldots,N}$
8:  for each timestep t do
9:     Observe state $\text{Observe state }s_{t}.$
10:     $\textbf{P}^{max}\leftarrow\text{Maximum power consumption over }H$
11:     $\{\textbf{T}^{sp}\}_{i=1,\ldots,N}\leftarrow\text{Temperature setpoints over }H$
12:     Run the constraint conversion Algorithm 2
13:     if DO ADMM then
14:        Initialize $\text{Initialize }\boldsymbol{\Delta},\bar{\textbf{u}},\{\boldsymbol{\lambda}\}_{i=1,\ldots,N}$
15:        while has not converge do
16:           $\{\bar{\textbf{u}}\}_{i=1,\ldots,N}\leftarrow Agg(\textbf{P}^{tot})$
17:           ($(\{\textbf{u}\}_{i=1\ldots N},\{\boldsymbol{\Delta}\}_{i=1\ldots N})\leftarrow LC_{i}(\bar{\textbf{u}}_{i},\boldsymbol{\lambda}_{i})$
18:           $\boldsymbol{\lambda}_{i}\leftarrow\boldsymbol{\lambda}_{i}+\rho(\bar{\textbf{u}}^{k+1}_{i}-\textbf{u}^{k}_{i})$
19:        end while
20:     end if
21:     $u_{t,i}\leftarrow ST_{i}(T^{sp}_{i},\delta_{t,i})$
22:     $Env.step(u_{t,i})$
23:     update $\text{update }\{\mathcal{D}\}_{i=1,\ldots,N}$
24:     if enough timesteps then
25:        Update $\text{Update }LC_{i}\text{ on }\mathcal{D}_{i}$
26:     end if
27:  end for

### v2

@@
-A Distributed ADMM-based Deep Learning Approach for Thermal Control in Multi-Zone Buildings
+A Distributed ADMM-based Deep Learning Approach for Thermal Control in Multi-Zone Buildings under Demand Response Events
 Abstract
-The surge in electricity use, coupled with the dependency on intermittent renewable energy sources, poses significant hurdles to effectively managing power grids, particularly during times of peak demand.
-Demand Response programs and energy conservation measures are essential to operate energy grids while ensuring a responsible use of our resources This research combines distributed optimization using ADMM with Deep Learning models to plan indoor temperature setpoints effectively.
+The increasing electricity use and reliance on intermittent renewable energy sources challenge power grid management during peak demand, making Demand Response programs and energy conservation measures essential.
+This research combines distributed optimization using ADMM with deep learning models to plan indoor temperature setpoints effectively.
 A two-layer hierarchical structure is used, with a central building coordinator at the upper layer and local controllers at the thermal zone layer.
@@
 Local controllers can modify the temperature setpoints to meet the local power targets.
-The resulting control algorithm, called Distributed Planning Networks, is designed to be both adaptable and scalable to many types of buildings, tackling two of the main challenges in the development of such systems.
+While most algorithms are either centralized or require prior knowledge about the building’s structure, our approach is distributed and fully data-driven.
+The proposed algorithm, called Distributed Planning Networks, is designed to be both adaptable and scalable to many types of buildings, tackling two of the main challenges in the development of such systems.
 The proposed approach is tested on an 18-zone building modeled in EnergyPlus.
@@
 Deep Learning, Distributed Optimization, ADMM, Demand Response, HVAC, Buildings
-This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
 I Introduction
@@
 More than half of the building’s energy is used to maintain the occupant’s comfort, mainly through Heating, Ventilation and Air Conditioning (HVAC) systems [1].
-Buildings are powerful pillars for integrating energy efficiency and Demand Response (DR) measures to reduce electricity consumption, especially during unstable power supply and peak periods.
+Buildings are crucial for integrating energy efficiency and Demand Response (DR) measures to reduce electricity consumption, especially during unstable power supply and peak periods.
 Many paths are investigated to benefit grid management by leveraging the building’s power consumption.
@@
 Optimization problems solved using MPC or RL are often formulated as centralized problems.
-In centralized problems, a single entity computes the solution to the problem and takes all the decisions.
+In centralized problems, a single entity computes the solution to the problem and makes all the decisions.
 Centralized problem formulations for BEMS are further described in [16].
@@
 In fully distributed architectures, LCs directly communicate with each other to coordinate their actions.
-The choice of architecture is problem dependent, as each architecture has pros and cons and highly depends on the communication infrastructure [18, 11].
+The choice of architecture is problem-dependent, as each architecture has pros and cons and highly depends on the communication infrastructure [18, 11].
 Distributed architectures have been increasingly used along with the Alternating Direction Method of Multipliers (ADMM) to tackle energy management challenges at the scale of the building.
@@
 The use of a non-convex ADMM allows to consider models more sophisticated than linear ones.
-In particular, models relying on Neural Networks proved to be extremely accurate to forecast the energy consumption in buildings [31, 23].
-Recurrent State Space Models (RSSM) also proved to be extremely accurate to model stochastic environments and are yet to be applied to building energy forecasting.
+In particular, models relying on neural networks proved to be extremely accurate in forecasting the energy consumption in buildings [31, 23].
+Recurrent State Space Models (RSSM) also proved to be extremely accurate in modeling stochastic environments and are yet to be applied to building energy forecasting.
 Such models, along with planning algorithms provide state-of-the-art results in many benchmark control tasks [32, 33].
@@
 While optimizing the HVAC system’s operation to efficiently track a setpoint is important, it is crucial to note that for a given temperature setpoint, there exists an inherent physical constraint on the extent of energy reduction achievable solely through the optimization of HVAC system operations.
-Further energy reduction may only be achieved by altering the setpoint itself.
+Further energy reduction may only be achieved by modifying the setpoint itself.
 This study thus uses temperature setpoints as actuators for power consumption regulation.
@@
 First, we formulate the temperature control problem as a centralized problem with a constraint on the building’s total power consumption.
-Then, we convert this problem into a non-convex consensus problem that we decompose using the ADMM algorithm.
+Then, we convert this problem into a non-convex sharing problem that we decompose using the ADMM algorithm.
 We provide convergence proof of the algorithm for this non-convex setting, which is key to accurately account for the non-linear dynamics of buildings.
 From this formulation, we derive a surrogate problem that can be efficiently solved in practice.
-We test both RSSM and SSM to model the building’s zone.
+We test both State Space Models (SSM) and RSSM to model the building’s zone.
 The models are used to plan for the temperature setpoints that best match the comfort requirements of zones while enforcing a limit on the building’s power usage.
@@
   1. 1.
-    This work proposes an efficient algorithm to adjust a building’s indoor temperature based on available energy resources.
-The approach contributes to energy conservation while minimally affecting occupant comfort, making it a valuable asset for DR.
+    We combine non-convex ADMM and fully data-driven deep-learning models to develop a distributed control algorithm for HVAC systems.
+A convergence analysis of the ADMM algorithm in this non-convex setting is presented.
+To the best of our knowledge, this is the first work combining non-linear deep-learning models and ADMM.
   2. 2.
-    We combine non-convex ADMM and advanced, fully data-driven, Deep Learning models to provide a distributed control for HVAC systems.
-To the best of our knowledge, this is the first work combining non-linear Deep Learning models and ADMM.
+    We propose a hard constraint of the maximum HVAC power consumption to prioritize energy savings during peak periods, while the use of comfort intervals instead of fixed temperature setpoints guarantees a minimum level of comfort for the users.
   3. 3.
-    While the global constraint is on the power consumption of the building, we propose to use temperature setpoints of the zones as actuators.
-Therefore, no assumption is made on the type of HVAC system, which makes our approach easily adaptable from one building to another.
-  4. 4.
-    The building is modeled using a fully data-driven approach, where no prior knowledge about the building’s blueprint is required.
-The RSSM allows one to plan in a stochastic environment and handle uncertainties in a robust MPC framework.
-The SSM can be used in computationally efficient shooting controllers.
-  5. 5.
-    We propose a convergence analysis for our non-convex ADMM algorithm. A case study shows the good convergence properties of our algorithm in practice.
-II Problem formulation
-In this study, a building composed of $N$ thermal zones is considered.
-Each zone has a temperature control system, i.e.
-a device like a PID or ON/OFF controller able to track a local temperature setpoint.
-Specifically, we consider a discretized setting, where occupants choose temperature setpoints $T^{sp}_{t,i}$ for each zone, that are optimal for their comfort.
+    We derive a deterministic and a stochastic version of our algorithm.
+The performances of both approaches are evaluated on a multi-zone building during DR events.
+The results demonstrate the control system’s capability to lower power consumption during peak periods.
+The algorithm consistently converges to a solution with each iteration.
+Additionally, the distributed architecture of the control system improves its scalability as the building size increases.
+In section II, we introduce the control problem and derive a distributed algorithm to solve it.
+In section III, we present the deep learning models used to model the thermal zones and the implementation of the proposed distributed control algorithm.
+In section IV, we present experimental results on the prediction accuracy of the models and the control performances of the algorithm.
+Finally, we discuss the results and some limitations paving the way for future works in section V.
+II Description of the Control Problem
+II-A Problem Formulation
+Consider a building composed of $N$ thermal zones equipped with independent HVAC systems.
+In each zone, occupants may choose a temperature setpoint $T^{sp}_{t,i}$ that is optimal for their comfort, where $t$ is a time index and $i$ a zone index.
 These setpoints may vary in time but are known beforehand.
-At time $t$ for the zone $i$, $u_{t,i}$ represents the power consumption of a Heating or Cooling (HC) system.
-The power consumption is closely related to the temperature setpoints [35].
-One can act on the setpoints to control the building’s HC power consumption below a maximum value.
-To this purpose, the global power constraint must be distributed to local temperature controllers acting on each zone, and translated to temperature setpoints.
-The proposed control system modifies the occupants’ setpoints $T^{sp}_{t,i}$ by $\delta_{t,i}$, so that the setpoint tracked by the HC system is $T^{sp}_{t,i}+\delta_{t,i}$ at time $t$ is zone $i$.
+To track their setpoint, each HVAC system consumes a given amount of power $u_{t,i}$.
+The HVAC power consumption is directly impacted by the temperature setpoints [35] and we propose to use the setpoints as actuators to control the HVAC consumption.
+Specifically, we aim at changing the occupants’ preferred setpoints $T^{sp}_{t,i}$ by $\delta_{t,i}$, so that the HVAC system tracks the setpoint $T^{sp}_{t,i}+\delta_{t,i}$.
 Let
@@
 $\textbf{u}=[u_{1,1},\ldots,u_{1,N},\ldots,u_{H,1},\ldots u_{H,N}]^{T}.$
-The optimal setpoint changes $\boldsymbol{\Delta}$ are solutions to the following optimization problem:
-{mini!}
-[…]
+The purpose of the control system is to apply a constraint on the entire building’s HVAC maximum power consumption.
+We define the optimal setpoint changes $\boldsymbol{\Delta}$ to satisfy this constraint as solutions to the following optimization problem:
+$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\|\boldsymbol{\Delta}\|^{2}_{2}$ (1)
+subject to $\boldsymbol{s}_{t+1,i}=f_{i}\left(\boldsymbol{s}_{t,i},\delta_{t,i},\boldsymbol{s}_{y,j\in\mathcal{N}_{i}},\delta_{t,j\in\mathcal{N}_{i}}\right),\forall t,i$ (2)
+$m_{i}\leq\delta_{t,i}\leq M_{i},\forall t,i$ (3)
+$A\textbf{u}\leq\textbf{P}^{max}$ (4)
 where
   • $f_{i}$ represents the dynamics of zone $i$,
-  • $s_{t,i}$ is the state of zone $i$.
-The state includes, in particular, $T^{sp}_{t,i}$, the air temperature, and the HC power consumption.
-The state is partially observable as it may also include the temperature of the walls or the thermal capacity of the zone.
+  • $\mathbb{s}_{t,i}$ is the state of zone $i$.
+The state includes, in particular, $T^{sp}_{t,i}$, the air temperature, and the HVAC power consumption.
+Note that the state is partially observable as it may also include the temperature of the walls or the thermal capacity of the zone.
   • $\mathcal{N}_{i}$ is the set of neighbors of zone $i$,
@@
     $A=\begin{bmatrix}1&1&0&0\\ 0&0&1&1\end{bmatrix}.$
-    The coefficient $A_{1,1}$ selects the power consumption of zone $1$ at the first time step.
-  • $\textbf{P}^{max}=[P^{max}_{1},\ldots,P^{max}_{H}]$ is a vector containing the maximum power for each timestep. The vector inequality in eq. (II) is component-wise.
-  • $M_{i}$ is the maximum setpoint change defined by the occupants in zone $i$.
-To solve this problem efficiently, we reformulate it as a sharing problem [29] and apply ADMM.
-First, the inequality constraint (II) is relaxed using a new parameter $\textbf{P}^{tot}$ such that $\textbf{P}^{tot}\leq\textbf{P}^{max}$.
+  • $\textbf{P}^{max}=[P^{max}_{1},\ldots,P^{max}_{H}]$ is a vector containing the maximum power for each timestep. The vector inequality in Eq. (4) is component-wise.
+  • $m_{i}$ and $M_{i}$ are the minimum and maximum setpoint changes defined by the occupants in zone $i$.
+This formulation prioritizes power savings over temperature comfort as long as the temperature setpoint stays in the comfort interval $\left[T^{sp}_{t,i}-m_{i};T^{sp}_{t,i}+M_{i}\right]$ defined by the user.
+To solve Problem (1) efficiently, we reformulate it as a sharing problem [29] and apply ADMM to solve it.
+First, the inequality constraint (4) is relaxed using a new parameter $\textbf{P}^{tot}$ such that $\textbf{P}^{tot}\leq\textbf{P}^{max}$.
 This ensures the problem is feasible and removes a coupling constraint.
@@
 In many buildings such as hotels, offices, or residential buildings, the temperature ranges in the zones are similar, thus limiting the heat exchanges.
-We will assess this assumption experimentally in section V to show that it is reasonable if the zones are well insulated.
+We will assess this assumption experimentally in section IV to show that it is reasonable if the zones are well insulated.
 Such an assumption is also discussed in [22, 36].
 The optimization problem becomes:
-{mini!}
-[…]
-The HC power consumption is now driven towards the parameters $\textbf{P}^{tot}=[P_{1}^{tot},\ldots,P_{H}^{tot}]$.
-The choice of $\textbf{P}^{tot}$ is important and will be further discussed in section V.
-Let us define a function $\phi_{i}$ that maps the state of a zone $i$ to its power usage:
-$u_{t,i}=\phi_{i}(s_{t,i}).$
-Given the setpoints and the first state, future states are computed by unrolling the dynamics:
-$s_{1,i}$ $=q_{i}(s_{0,i},\delta_{0,i})$
-$s_{2,i}$ $=q_{i}(s_{1,i},\delta_{1,i})$
-$=q_{i}(q_{i}(s_{0,i},\delta_{0,i}),\delta_{1,i})$
-$\ldots$
-$s_{H,i}$ $=q_{i}(s_{H-1,i},\delta_{H,i})$
-$=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{H,i})$
-To ease the notations, we define the following functions, for all $i$ and $t$:
-$Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})=q_{i}(q_{i}(\ldots q_{i}(s_{0,i},\delta_{0,i})),\delta_{t,i})$ (1)
+$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\|\boldsymbol{\Delta}\|^{2}_{2}+||A\textbf{u}-\textbf{P}^{tot}||^{2}_{2}$ (5)
+subject to $\boldsymbol{s}_{t+1,i}=q_{i}\left(\boldsymbol{s}_{t,i},\delta_{t,i}\right),\quad\forall t,i$ (6)
+$m_{i}\leq\delta_{t,i}\leq M_{i},\forall t,i$ (7)
+Note that we do not use weights to balance the two terms of the objective.
+In practice, the temperature and power values should be normalized for the two terms to have the same order of magnitude.
+The HVAC power consumption is now driven towards the parameters $\textbf{P}^{tot}=[P_{1}^{tot},\ldots,P_{H}^{tot}]$.
+The choice of $\textbf{P}^{tot}$ is important and will be further discussed in section III-B3.
+The powers u are components of the state s and the constraint Eq.
+(6) can be made implicit by unrolling the dynamics $q_{i}$ from the initial state using the setpoint schedules.
+To this purpose, let $\phi_{i}$ map the state of a zone $i$ to its power usage $u_{t,i}=\phi_{i}(\mathbb{s}_{t,i})$, and consider the following functions for all $i$ and $t$:
+$Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})=q_{i}(q_{i}(\ldots q_{i}(\mathbb{s}_{0,i},\delta_{0,i})),\delta_{t,i})$ (8)
 $g(\boldsymbol{\Delta}_{i})=\left\lVert\boldsymbol{\Delta}_{i}\right\rVert^{2}_{2}$ (9)
-$\ell(\boldsymbol{\Delta})=\sum_{t=1}^{H}\left(\sum_{i=1}^{N}\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})-\textbf{P}^{tot}_{t}\right)^{2}$ (3)
-The index to the function $Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i})$ indicates that it represents the state at time $t$ of zone $i$ computed from $s_{0,i}$.
-Only the first $t-1$ elements of $\boldsymbol{\Delta}_{i}$ are needed to compute $s_{t,i}$, but the entire vector is passed for notational simplicity.
-The power usage becomes:
-$u_{t,i}=\phi_{i}(Q_{t,i}(s_{0,i},\boldsymbol{\Delta}_{i}))$ (4)
-[…]
-This problem is a sharing problem as described in [29].
+$\ell(\boldsymbol{\Delta})=\sum_{t=1}^{H}\left(\sum_{i=1}^{N}\phi_{i}(Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})-\textbf{P}^{tot}_{t}\right)^{2}$ (10)
+The index to the function $Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})$ indicates that it represents the state at time $t$ of zone $i$ computed from $\mathbb{s}_{0,i}$.
+Only the first $t$ elements of $\boldsymbol{\Delta}_{i}$ are needed to compute $\mathbb{s}_{t,i}$, but the entire vector is passed for notational simplicity.
+The power usage of a zone is thus
+$u_{t,i}=\phi_{i}(Q_{t,i}(\mathbb{s}_{0,i},\boldsymbol{\Delta}_{i})).$ (11)
+Using these notations, Problem (5) may be reformulated as a sharing problem
+$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\sum_{i=1}^{N}g(\boldsymbol{\Delta}_{i})+\ell(\boldsymbol{\Delta})$ (12)
+subject to $m_{i}\leq\delta_{t,i}\leq M_{i}\quad\forall t,i.$ (13)
 Each zone has a local objective represented by $g$ and a global objective represented by $\ell$.
 The total power available is a common resource that must be shared across all zones to satisfy the local comfort requirements.
-[…]
+II-B Distributed Sharing Algorithm
+In this section, we derive a distributed algorithm based on ADMM to solve the Problem (12).
+Note that no particular assumption is made on the dynamics $q_{i}$ of the zones and the problem is thus non-convex.
+First, we introduce duplicated variables $\bar{\boldsymbol{\Delta}}_{1},\ldots,\bar{\boldsymbol{\Delta}}_{N}$ and form an equivalent problem with $N$ vector linear equality constraints:
+$\underset{\boldsymbol{\Delta}}{\text{minimize}}\quad$ $\sum_{i=1}^{N}g(\boldsymbol{\Delta}_{i})+\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{\Delta}}_{i}\right)$ (14)
+subject to $\bar{\boldsymbol{\Delta}}_{i}=\boldsymbol{\Delta}_{i}\quad\forall i$ (15)
+$m_{i}\leq\delta_{t,i}\leq M_{i}\quad\forall t,i$ (16)
 The matrices $B_{i}\in\mathbb{R}^{NH\times H}$ for $i=1,\ldots,N$ are such that $\boldsymbol{\Delta}=\sum_{i=1}^{N}B_{i}\boldsymbol{\Delta}_{i}$.
@@
 11:  end while
-The sharing problem can be solved using Alg.
-1.
+The sharing problem can be solved using Algorithm 1.
 We distinguish two main sub-problems in this algorithm.
-The first one is the local controller problem (line 5) and the second one is the coordinator problem (line 7).
+The first one is the Local Controller (LC) problem (line 5) and the second one is the coordinator problem (line 7).
 As stated in the following theorem, under some Lipschitz regularity assumptions and the right choice of $\rho$, this algorithm is guaranteed to converge although the problem is not convex.
-We will assess the quality of the solutions experimentally in section V.
+We will assess the quality of the solutions experimentally in section IV.
+A proof of the theorem is given the Appendix A-A.
 Theorem II.1.
-The Alg. 1 converges to the set of stationary solutions of problem II :
+The Algorithm 1 converges to the set of stationary solutions of Problem (14):
 $\lim_{k\xrightarrow{}\infty}dist(\{\boldsymbol{\Delta}_{i}^{k}\},\{\bar{\boldsymbol{\Delta}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\};Z^{*})=0,$
 where $Z^{*}$ is the set of primal-dual stationary solutions of the problem.
-Evaluating the models of each dynamics $q_{i}$ is computationally expensive.
-To leverage the benefits of the decomposition, the models’ evaluation should be placed in the for loop and computed in parallel.
-In Alg.
-1 the models’ evaluation is in the $\ell$ function, in the coordinator problem.
-We can explicit the relationship between the setpoints and the related power consumption (4) as a constraint in the LC problem.
-{mini!}
-[…]
-and consider the surrogate coordinator problem:
-{mini!}
-[3] ¯u~ℓ(¯u) + ∑_i=1^N λ_i^T (¯u_i - u_i) + ρ—— ¯u_i - u_i ——_2^2
-where ~$\tilde{\ell}(\bar{\textbf{u}})=\left\lVert A\bar{\textbf{u}}-\bar{P}^{tot}\right\rVert$.
-This formulation may be derived from the same decomposition steps as the ones above, but using the following equation rather than eq.
-(4):
+Querying the zones’ dynamics $q_{i}$ is computationally expensive.
+To leverage the benefits of the decomposition, the dynamics’ evaluation should be placed in the for loop and computed in parallel.
+In Algorithm 1 the dynamics are called in the $\ell$ function, in the coordinator problem.
+It means that the local controllers are sending their temperature setpoint schedules to the coordinator which has to query all the zones’ dynamics to compute the associated HVAC consumption.
+Algorithm 1 formulation is convenient for a theoretical analysis of the problem.
+In practice, however, it is more efficient to compute the HVAC consumption at the LC level and exchange power consumption messages with the coordinator.
+We thus explicit the relationship between the setpoints and the related power consumption given by Eq.
+(11) as a constraint in the LC problem and derive a surrogate LC problem:
+$\underset{\boldsymbol{\Delta}_{i}}{\text{minimize}}\quad$ $g(\boldsymbol{\Delta}_{i})+\boldsymbol{\lambda}_{i}^{T}(\bar{\textbf{u}}_{i}-\textbf{u}_{i})+\rho||\bar{\textbf{u}}_{i}-\textbf{u}_{i}||_{2}^{2}$ (18)
+subject to $u_{t,i}=\phi(Q_{t,i}(\boldsymbol{s}_{0,i},\boldsymbol{\Delta}_{i}))$ (19)
+$\bar{u}_{t,i}=\phi(Q_{t,i}(\boldsymbol{s}_{0,i},\bar{\boldsymbol{\Delta}}_{i}))$ (20)
+$m_{i}\leq\delta_{t,i}\leq M_{i}$ (21)
+as well as a surrogate coordinator problem:
+minimizeu¯~($\underset{\bar{\textbf{u}}}{\text{minimize}}\quad\tilde{\ell}(\bar{\textbf{u}})+\sum_{i=1}^{N}\boldsymbol{\lambda}_{i}^{T}(\bar{\textbf{u}}_{i}-\textbf{u}_{i})+\rho||\bar{\textbf{u}}_{i}-\textbf{u}_{i}||_{2}^{2}$ (22)
+where ~$\tilde{\ell}(\bar{\textbf{u}})=\left\lVert A\bar{\textbf{u}}-P^{tot}\right\rVert$.
+In our implementation, these surrogate problems are used in place on the LC and coordinator problem in Algorithm 1.
+The resulting control architecture is summarized in Fig.
+1.
+Note that the surrogate formulation may be derived from the same decomposition steps as the ones presented in Section II-A, but using the following equation rather than Eq. (11):
 $\delta_{t,i}=\psi(Q_{t,i}(s_{0,i},u_{i})),$ (23)
@@
 For instance, a proportional controller that operates with a zone’s temperature around the setpoint or an ON/OFF controller used with Pulse Width Modulation would give good properties to $\psi$.
-Instead of enumerating such conditions, that are hard to verify in practice, we choose to provide a theoretical analysis for the problem formulation (4).
-This formulation is more natural and indicates the form of the ADMM decomposition for the surrogate problem.
-As shown in Section V, the surrogate formulation yields efficient computation and shows excellent convergence properties experimentally.
+Instead of enumerating such conditions, that are hard to verify in practice, we provide a theoretical analysis for the Problem (12).
+This formulation arises naturally from Problem (1) and indicates the form of the ADMM decomposition for the surrogate problem.
+As shown in Section IV, the surrogate formulation yields efficient computation and shows excellent convergence properties experimentally.
 In addition, it provides a good intuition about the underlying mechanism: the coordinator (i.e.
-problem (II.1)) is choosing the amount of power for each room and each timestep, to ensure that the power constraint is always satisfied.
+Problem (22)) is choosing the amount of power for each room and each timestep, to ensure that the power constraint is always satisfied.
 The LCs (i.e.
-problems (II.1)) must then find the temperature setpoints to track this power target while preserving the users’ comfort.
-Even if multiple setpoints may satisfy the constraints II.1 and II.1, the lowest setpoint changes will be chosen because of the term $g(\Delta_{i})$ is the objective function.
-The control architecture is summarized in Fig. 1.
-Figure 1: Two layer hierarchical control architecture.
+Problems (18)) must then find the temperature setpoints to track this power target while preserving the users’ comfort.
+Even if multiple setpoints may satisfy the constraints Eq.
+(19) and (20), the lowest setpoint changes will be chosen because of the term $g(\Delta_{i})$ is the objective function.
+Figure 1: Two layers hierarchical control architecture.
 The LCs observe the state of their zone and plan for the best setpoints while communicating with the coordinator to enforce a maximum constraint on power usage.
 Setpoints are sent to the existing setpoint trackers that act on each zone.
-III Local Controllers and Zone Modelling
+III Distributed Planning Networks
+In this section, we present the methods used to solve the optimization problems of Algorithm 1.
+In practice, the dynamics $q_{i}$ of the zones are unknown.
+In this study, we consider deep learning models to learn the dynamics from data.
+Both a deterministic and a stochastic approach are used, as described in the first part of this section.
+The models are used for planning to solve the LC problem (18), yielding the Distributed Planning Networks (DPN) algorithm.
+In the second part this section, we present a deterministic and a stochastic version of the DPN algorithm.
+III-A Deep Learning Models for Thermal Zones
 Figure 2: Distributed Planning Network - For each iteration of the ADMM algorithm, the coordinator sends power constraints to the LCs.
@@
 Once the ADMM has converged, only the first action of the horizon is applied to the zones.
-To successfully plan for the best setpoint changes without implementing any actions, LCs must be able to predict the evolution of the system.
-In this section, we present two methods for the modeling of a thermal zone: Recurrent State Space Models (RSSM) and State Space Models (SSM).
-III-A Reccurent State Space Model
+To plan for the optimal setpoint changes, LCs must be able to predict the evolution of the thermal zones’ state.
+In this section, we present two types of models for the zones: State Space Models (SSM) and Recurrent State Space Models (RSSM).
+These models use inputs that include the observation of the zone $\textbf{o}_{t,i}$, actions $\textbf{a}_{t,i}$ and disturbances $\textbf{d}_{t}$ like weather and calendar information, to predict the next state in the observable space.
+To not overload the prediction models, the disturbances are not predicted by the models but given as inputs, as pictured in Fig.
+3.
+One advantage of this approach is that a single weather prediction system is required for the entire building, and the zones’ model can focus on predicting only the zones’ states.
+The zones are partially observable environments and making predictions in a latent space helps to cope with partial observability.
+III-A1 State Space Models
+The SSM has an encoder-decoder structure and the transition function is a recurrent cell that takes as input the actions and disturbances.
+The latent state $s_{t}$ is carried by the hidden state of the recurrent cell:
+$\mathbb{s}_{t+1}$ $=f_{\boldsymbol{\theta}}(\mathbb{s}_{t},\mathbb{a}_{t},\mathbb{d}_{t})$ (24)
+$\mathbb{o}_{t+1}$ $=\text{dec}_{\boldsymbol{\theta}}(\mathbb{s}_{t+1})$ (25)
+$\mathbb{s}_{0}$ $=\text{enc}_{\boldsymbol{\theta}}(\mathbb{o}_{-1},\ldots\mathbb{o}_{-n})$ (26)
+where enc is an recurrent neural network encoding lags of observations, dec is a neural network decoding the latent state and $f$ is a recurrent cell used as the transition function.
+Prior works such as [23] have used similar models for representing buildings’ dynamics.
+The model is trained to minimize the mean squared error between the predicted and actual powers on the entire prediction horizon.
+III-A2 Reccurent State Space Model
 The RSSM is a stochastic model that allows forecasting of the power usage of each zone under uncertainties brought by many unobservable disturbances.
@@
   1. 1.
-    A representation model (encoder), which maps the observations $o_{t}$ to the latent state $s_{t}$ of the system.
-The use of an encoder is important because the system is partially observed.
-One can include lags of power and temperature in the observation to build a better representation of the state and cope with partial observability.
+    A representation model (encoder), which maps the observations $\mathbb{o}_{t}$ to the latent state $\mathbb{s}_{t}$ of the system.
   2. 2.
     A transition model, that predicts the evolution of the latent space.
-Note that this transition model is composed of two parts: (1) a deterministic state model which predicts the evolution of the deterministic part of the state $h_{t}=f(h_{t-1},s_{t-1},\delta_{t-1})$ and (2) a stochastic state model which samples the next latent state according to $h_{t}$ : $s_{t}\sim p(s_{t}|h_{t})$.
+Note that this transition model is composed of two parts: (1) a deterministic state model which predicts the evolution of the deterministic part of the state $\boldsymbol{h}_{t}=f(\mathbb{h}_{t-1},\mathbb{s}_{t-1},\delta_{t},\mathbb{d}_{t})$ and (2) a stochastic state model which samples the next latent state $\mathbb{s}_{t}\sim p(\mathbb{s}_{t}|\mathbb{h}_{t})$.
 The latter part is important in our case as it accounts for the perturbations due to the occupancy and neighboring zones.
@@
 The objective is to maximize the variational bound as proposed in [32].
-The transition model allows to predict one timestep ahead in the latent space.
-Multiple steps prediction is achieved by applying the model recursively.
-This is convenient as the user may want to change the prediction horizon without having to retrain the model.
-The latent space embeds the inner state of the room, mainly composed of the power usage and the room’s temperature.
-This state is influenced by exogenous perturbations such as weather and the temperature setpoint.
-To not overload the transition model, we do not include these disturbances in the latent space as it is usually done with this type of model.
-Instead, we feed the transition model with the previous latent space as well as forecasts of the disturbances, as pictured in Fig.
-3.
-The advantages of this approach are that a single weather prediction system is required for the entire building.
-Since the RSSM of each zone does not have to predict the complex evolution of the weather on top of the evolution of the room’s state, it can provide better predictions using a smaller network model.
-In addition, some disturbances such as temperature setpoint schedules may be known beforehand and can be passed exactly to the transition model at each prediction step.
+The transition model predicts one timestep ahead in the latent space.
+Multiple-step prediction is achieved by applying the model recursively.
 Figure 3: Multi-step ahead power predictions with weather forecasts
-III-B State Space Models
-The State Space Model (SSM) has the same structure as the RSSM, but it does not include the stochastic part.
-The transition model is a GRU cell that takes as input the disturbances (i.e.
-the weather and the setpoints).
-The latent state $s_{t}$ is carried by the hidden state of the GRU cell.
-The model is trained to minimize the mean squared error between the predicted and actual powers.
-IV Distributed Planning Networks
-In this section, we present how the models of each LC are combined with the coordinator using ADMM.
-The proposed algorithm is called Distributed Planning Networks as it embeds the idea of the Planning Networks in a distributed architecture.
-Fig.
-2 and Alg.
-3 summarize the approach.
-The coordinator problem is a convex penalized least square problem. It can be solved easily with existing solvers. The challenge is more on the side of the LC problem (II.1).
-To efficiently find a solution to the non-convex LC problem, we propose two approaches.
-The first one is a random search for the best temperature setpoints.
-The second one is a shooting method, where we differentiate through the prediction model and use the gradient to guide the search.
+III-B Distributed Planning Networks Algorithms
+In this section, we present practical methods for solving the LC and coordinator optimization problems introduced in Algorithm 1.
+The coordinator problem is a convex penalized least square problem.
+It can be solved easily with existing solvers.
+The challenge is more on the side of the LC problem (18).
+For this problem, the zones’ models are used to plan for the best setpoint.
+We propose two approaches: the first one is a shooting method, where we differentiate through the prediction model and use the gradient to guide the search; the second one is a random search for the best temperature setpoints.
 Once the best setpoint changes have been found for the upcoming horizon, only the setpoint changes of the first timestep are implemented.
 The process is then repeated to allow for re-planning and correction of the control.
-IV-A Planning by Random Search with RSSM
-The random search is conducted as follows:
+The proposed algorithm is called Distributed Planning Networks (DPN) as it embeds Planning Networks in a distributed architecture. Fig. 2 and Algorithm 2 summarize the approach.
+III-B1 Deterministic DPN (DDPN)
+The SSM is fully differentiable.
+One can formulate the Problem (18) as an initial value problem and use single shooting to minimize the running cost given by the objective function.
+The equality constraints representing the dynamics are implicitly satisfied by the prediction model.
+To handle the inequality constraint (21), we use projected gradient descent to keep the setpoint changes within the desired interval.
+This approach assumes continuous values for the setpoints.
+In practice, we round the values at the resolution of the setpoint tracker of the zone afterward.
+A stopping criterion on the number of decimals changing in $\Delta_{i}$ at each iteration is used to make the approach computationally efficient.
+III-B2 Stochastic DPN (SDPN)
+The random search is conducted as follow:
   1. 1.
@@
 For each $\boldsymbol{\Delta}_{i}$, the RSSM outputs a distribution for the trajectory of states.
-For one choice of $\boldsymbol{\Delta}_{i}$, multiple states trajectories may be sampled from the distribution to account for uncertainties.
+For one choice of $\boldsymbol{\Delta}_{i}$, multiple state trajectories are sampled from the distribution to account for uncertainties.
+The trajectory with the maximum power consumption is kept to optimize on the worst-case scenario.
   3. 3.
-    Select $\boldsymbol{\Delta}_{i}$ providing the lowest score as defined by (II.1).
+    Select $\boldsymbol{\Delta}_{i}$ providing the lowest score as defined by (18).
 While using ADMM, the steps presented above need to be repeated at each iteration.
@@
 To limit the computational time, the trajectories (i.e.
-the two first steps) are pre-computed and stored before the start of ADMM.
+the two first steps) are pre-computed and stored before the start of the ADMM iterations.
 Note that these two steps can be performed in parallel for each zone.
 At each ADMM iteration, the score of each trajectory is then computed from the pre-computed trajectory with respect to the current power targets $\bar{\textbf{u}}_{i}$.
-IV-B Planing by Shooting with SSM
-The SSM is fully differentiable.
-One can formulate the problem (II.1) as an Initial Value Problem and use single shooting to minimize the running cost given by the objective function.
-The equality constraints representing the dynamics are implicitly satisfied by the prediction model.
-To handle the inequality constraint (II.1), we use projected gradient descent to keep the setpoint changes within the desired interval.
-This approach assumes continuous values for the setpoints.
-In practice, we round the values at the resolution of the setpoint tracker of the zone afterwards.
-A stopping criterion on the number of decimals changing in $\Delta_{i}$ at each iteration is used to make the approach computationally efficient.
-V Results
-To test the proposed approach, we use the low-rise apartment building, one of the Department Of Energy (DOE) residential reference buildings.
-The building is modeled in EnergyPlus, a software that provides thermal simulations for buildings with state-of-the-art accuracy.
-The building is composed of three floors of 6 apartments each.
-Apartments are equipped with independent HVAC systems, temperature setpoints schedules, occupancy schedules, and appliance usage schedules.
-Simulations are run using Typical Meteorological Year (TMY) files coming from the DOE climate zone 7A, representing the weather in the North-East of the United States of America and South-East of Canada.
-In each apartment, the user setpoint $T_{i}^{sp}$ is kept constant, but may vary between $19^{\circ}C$ and $22^{\circ}C$ from one apartment to the other.
-For the purpose of this study, we are considering heating periods, and hence we modify the default HVAC systems to electric heating-only systems with an efficiency of one.
-The EnergyPlus timestep was set to 15 minutes.
-The prediction horizon is set to one or two hours and the ADMM timestep is 30 minutes, meaning we can re-plan to change the setpoints every 30 minutes.
-V-A Models accuracy
-zone RSSM SSM zone RSSM SSM
-1h ahead forecast 1 185 78 10 125 89
-2 129 79 11 158 140
-3 181 88 12 127 71
-4 136 67 13 177 92
-5 122 55 14 134 94
-6 130 68 15 164 105
-7 137 94 16 152 69
-8 144 116 17 128 57
-9 159 107 18 134 84
-zone RSSM SSM zone RSSM SSM
-2h ahead forecast 1 184 80 10 98 82
-2 120 71 11 144 133
-3 182 88 12 109 72
-4 134 63 13 173 103
-5 104 51 14 112 96
-6 127 66 15 159 110
-7 127 105 16 122 71
-8 131 121 17 105 79
-9 152 108 18 109 102
-TABLE I: Models scores (MAE, Watt) for one hour and two hours ahead predictions during a three weeks test period.
-For the entire building’s power forecast, the RSSM has a MAPE of $5.81\%$ for one hour ahead and $6.20\%$ for two hours ahead.
-The SSM has a MAPE of $3.72\%$ and $4.58\%$ respectively.
-[…]
-First, we test the accuracy of the prediction models.
-Both RSSM and SSM are trained offline for each zone of the building.
-The training data are generated by running simulations with random temperature setpoints changed every period of random duration, ranging from 1h to 24h.
-The observation of a zone contains the zone’s current temperature as well as the power consumption of each timestep for the last hour.
-The output of the representation model is the power consumption of the next timestep.
-At each timestep, the zone’s observation is completed with disturbances inputs.
-Disturbances include the temperature setpoint and the temperature setpoint changes that are known beforehand.
-As one may observe in Fig.
-4, the changes in the temperature setpoints cause steep variations in the power.
-To accurately predict these changes, a key feature added is the difference between setpoint changes from one timestep to the next i.e.
-$\delta_{t,i}-\delta_{t-1,i}$.
-The disturbances also include the weather, composed of the outdoor temperature and solar radiation.
-For simplicity, we did not build a weather forecast model.
-Instead, we used perfect weather information and added random noise to account for prediction errors.
-Table I shows the Mean Absolute Error (MAE) of each zone for one-hour and two-hours ahead forecasts.
-The SSM performs consistently better than the RSSM, but both methods give accurate predictions.
-For the entire building’s power forecast, the Mean Average Percentage Errors (MAPE) of the SSM are $3.72\%$ and $4.58\%$ for one-hour and two-hours ahead predictions.
-The MAPE of the RSSM are $5.81\%$ and $6.20\%$ respectively.
-These predictions were made without observing the states of the neighboring zones, as stated in the assumption of the model used in (II).
-V-B Distributed Planning Networks and Power Control
-To test the Distributed Planning Networks algorithm, we propose to implement it for peak load reduction.
-We consider a Demand Response (DR) program where the utility sends an early notification to the building’s operator to reduce the power demand during peak demand events.
-We assume that the events occur between $5AM$ and $9AM$ every morning over a period of three weeks.
-We also assume that this reduction implies a limit on the heating demand that should not exceed $\textbf{P}^{max}_{t}=40kW$ for all timesteps $t$ during DR events.
-For each zone, we set $-2^{\circ}C\leq\boldsymbol{\Delta}_{t}\leq 0^{\circ}C$, with an action step of $0.25^{\circ}C$.
-Algorithm 2 Constraint conversion from $\textbf{P}^{max}$ to $\textbf{P}^{tot}$
-1:  Inputs : current state $\text{Inputs : current state }s_{t}\text{ and }\textbf{P}^{max}$
-2:  Prediction models $\text{Prediction models }\phi_{1},\ldots,\phi_{N}$
-3:  Parameters : $\text{Parameters : }\nu,\boldsymbol{\Delta}_{min},\boldsymbol{\Delta}_{max},H$
-4:  Initialize $\textbf{P}^{tot}\in\mathbb{R}^{H}$
-5:  for i=1, …, N do
-6:     $P_{normal,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{max}))$
-7:     $P_{min,i}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{min}))$
-8:  end for
-9:  $\textbf{P}_{normal}=\sum_{i=1}^{N}P_{normal,i},\text{ }\textbf{P}_{min}=\sum_{i=1}^{N}P_{min,i}$
-10:  if $\textbf{P}_{normal}\leq(1-\nu)\textbf{P}^{max}$ then
-11:     $\delta_{t,i}=0\text{ }\forall t,i$
-12:     DO ADMM = FALSE
-13:  else if $\textbf{P}_{min}>(1-\nu)\textbf{P}^{max}$ then
-14:     $\delta_{t,i}=-2\text{ }\forall t,i$
-15:     DO ADMM = FALSE
-16:  else
-17:     $\textbf{P}^{tot}=clip(\textbf{P}_{normal},(1-\nu)\textbf{P}^{max})$
-18:     DO ADMM = TRUE
-19:  end if
+III-B3 Maximum power constraint conversion
 As mentioned in Section II, translating the constraint’s parameter $\textbf{P}^{max}$ to $\textbf{P}^{tot}$ (defined in Eq.
-(II) ) is important as the building power consumption will be driven towards $\textbf{P}^{tot}$.
-To find the right $\textbf{P}^{tot}$, we propose to leverage the prediction model of each zone.
-We forecast the total power for setpoint changes of $0^{\circ}C$ and $-2^{\circ}C$ to find lower and upper bounds on the power.
-If the lower bound is below $\textbf{P}^{max}$, the optimal solution of the problem II is $\boldsymbol{\Delta}=0$ and no further computation is required.
-If the upper bound is above $\textbf{P}^{max}$, the problem (II) is unfeasible.
-In the latter case, depending on the energy pricing, one could either reduce as much as possible the temperatures to lower the consumption, or take no action not to alter the comfort since the DR requirement cannot be met.
-If $\textbf{P}^{max}$ falls between the upper and lower bounds, we proceed with the Distributed Planning Networks algorithm.
-We then set $\textbf{P}^{tot}=\textbf{P}^{max}$.
-Note that in practice, we use a slack parameter $\nu$ to mitigate the effects of prediction errors, and set $\textbf{P}^{tot}=(1-\nu)\textbf{P}^{max}$.
-This procedure is detailed in Alg.
-2.
-V-B1 Planning by shooting with SSM
-During the three week test period, the power consumption was successfully reduced below $\textbf{P}^{max}$ during each DR event.
-Fig.
-5 shows the power consumption during three days of the test period.
-Fig.
-6 shows the effect of the slack parameter $\nu$ on the heating power consumption.
-The more $\nu$ increases, the more conservative the control is but the higher the chance to satisfy the power constraint despite prediction errors.
-In our experiment, the power constraint was satisfied for the three choices $\nu=8\%$, $\nu=4\%$, and $\nu=0\%$.
-However, the maximum power monitored during DR event increased as the $\nu$ decreased.
-With $\nu=0\%$ the power reached $39.9kW$, compare to a maximum of around $38.5kW$ for $\nu=8\%$, making the power constraint really tight.
-The right choice of $\nu$ is therefore a trade-off between comfort and power usage.
-In Fig.
-5, we observe rebound peaks caused by the increase of temperature setpoints after the DR events.
-These peaks may be managed in several ways.
-For instance, as shown in Fig.
-7, one may extend the duration of the DR event to reduce the magnitude of the rebound peaks.
-Our algorithm may also be used for peak shaving by maintaining the power constraint throughout the day, as shown by the green curve in Fig.
-7.
-Table II shows the distribution of the temperature setpoints during active DR events.
-We observe that the effort is evenly spread among the zones.
-No zone had to take an extreme $-2^{\circ}C$ action and most of the actions are between $-0.5^{\circ}C$ and $\-1^{\circ}C$.
-Note that we did not enforce any type of fairness in the algorithm.
-We initialize the ADMM iterations with a setpoint change of $-1^{\circ}C$ and the same power target for every zone.
-We attribute this equal distribution of actions to the initialization.
-Differences arise because each zone has a different heating power requirement.
-Power consumption is influenced by many factors such as the orientation of the building, the position of the zone within the building, the presence of windows, and the occupancy.
-V-B2 Planing by random search with RSSM
-SSM zone mean $(^{\circ}C)$ None Low Medium High zone mean $(^{\circ}C)$ None Low Medium High
-1 -0.72 0 34.21 65.79 0 10 -0.61 0 52.63 47.37 0
-2 -0.56 0 65.79 34.21 0 11 -0.74 0 44.74 47.36 7.90
-3 -0.80 0 18.42 81.58 0 12 -0.78 0 26.31 71.05 2.64
-4 -0.73 0 34.21 65.79 0 13 -0.84 0 15.79 78.95 5.26
-5 -0.66 0 42.10 57.90 0 14 -0.61 0 55.26 44.74 0
-6 -0.78 0 21.05 78.95 0 15 -0.78 0 26.32 71.05 2.63
-7 -0.70 0 39.48 57.89 2.63 16 -0.82 0 18.42 78.95 2.63
-8 -0.66 0 50.0 44.73 5.27 17 -0.79 0 21.06 76.31 2.63
-9 -0.74 0 39.47 55.26 5.27 18 -0.73 0 26.32 71.05 2.63
-Model RSSM zone mean $(^{\circ}C)$ None Low Medium High zone mean $(^{\circ}C)$ None Low Medium High
-1 -1.14 8.12 13.51 24.32 54.05 10 -1.29 5.41 13.51 13.51 67.57
-2 -1.27 2.72 13.51 24.32 59.45 11 -1.32 2.70 8.11 18.92 70.27
-3 -1.24 2.71 10.81 29.72 56.76 12 -0.9 13.52 35.14 8.10 43.24
-4 -1.14 8.12 8.11 24.32 59.45 13 -1.17 8.12 21.62 13.51 56.75
-5 -1.35 0 18.92 16.22 64.86 14 -1.32 2.70 13.52 13.51 70.27
-6 -1.41 0 8.11 13.51 78.38 15 -0.78 0 51.35 16.22 32.43
-7 -1.23 0 16.21 24.32 59.46 16 -1.16 5.41 18.92 16.22 59.45
-8 -1.36 0 16.22 10.81 72.97 17 -1.17 0 24.33 21.62 54.05
-9 -1.41 0 8.10 16.22 75.68 18 -1.15 0 24.32 24.32 51.36
-TABLE II: Setpoint changes during active DR events (i.e.
-DR events where actions were required).
-The values correspond to the percentage of time spent with the following setpoint change : None := $\delta=0$, Low := $0.50<\delta\leq-$ 0.5, Medium := $-0.5<\delta\leq-1.0$, High := $-1.0<\delta\leq-1.75$.
-The results for the SSM are obtained with $\nu=4\%$.
-A random search is used along with the RSSM to plan for the best setpoints.
-We split the computational budget between exploring the action space (i.e.
-sampling trajectories for different $\boldsymbol{\Delta}_{i}$) and exploring the possible trajectories for each choice of $\boldsymbol{\Delta}_{i}$.
-At the beginning of each timestep, we randomly sample 100 different choices for $\boldsymbol{\Delta}_{i}$, and for each $\boldsymbol{\Delta}_{i}$, we sample 25 different trajectories.
-We then select the trajectory out of the 25 that yields the highest power consumption, to optimize based on the worst-case scenario.
-Fig.
-8 shows an example of different power consumptions sampled from a given state for a choice of $\boldsymbol{\Delta}_{i}$.
-Note that with this approach, we fix the slack parameter $\nu$ to zero because the optimization is already made based on a worst-case scenario.
-The experiment is repeated for 5 different seeds.
-The results are consistent across all the seeds.
-For simplicity, only one outcome is presented.
-Using this type of controller, the power is successfully controlled below $P^{max}$.
-The maximum power monitored during a DR event is $39.54\pm 0.56kW$  [^1] Mean and standard deviation computed on the different seeds., resulting in a tight power constraint.
-To build a better estimate of the worst-case scenario, one can increase the number of trajectories sampled for each $\Delta_{i}$.
-Alternatively, one may choose a strictly positive slack parameter to mitigate the impact of prediction errors on the satisfaction of the maximum power constraint.
-Fig.
-5 shows a three days example of power consumption.
-Note that the rebound peaks may also be managed by increasing the duration of DR events.
-As shown in Table II, the effort is spread across all the zones.
-As for the shooting method, the algorithm is initialized evenly by fixing the same power targets for every zone.
-As a result of the optimization based on the worst-case scenario, the temperature setpoint reductions are higher than with the shooting method.
-This results in a larger power reduction during DR events, as seen in Fig.
-5.
-One of the main limitations of this approach is the exploration of the action space.
-For a one-hour ahead planning, the action space is small and may be well explored by random search.
-However, to plan on a longer horizon, it is possible to reallocate the computational budgets to sampling more $\boldsymbol{\Delta}_{i}$ rather than different trajectories for a given $\boldsymbol{\Delta}_{i}$.
-The limiting factor is time, as the computation must be made within the $15$ minutes timestep.
-In our experiments, on a single CPU (3.5GHz), the sampling of all the $2500$ trajectories took around $5$ minutes.
-This computation is done in parallel for all the zones.
-Once the trajectories have been sampled, a hundred ADMM iterations can be performed within a minute.
-Figure 5: Comparison between the random search and the shooting control. Demand Response events occur between $5$AM and $9$AM every day. The slack is set to $\nu=4\%$ for the shooting control.
-Figure 6: Influence of the slack parameter $\nu$ on the power consumption for the shooting control.
-Figure 7: Influence of the duration of the DR event on the rebound peak. The black line represents $\textbf{P}^{max}$. Our algorithm is flexible and can also be used for peak shaving.
-Figure 8: Different power trajectories for a zone, sampled from a given state $\boldsymbol{\Delta}_{i}$ using a RSSM
-V-B3 Convergence Analysis
-Figure 9: Mean and standard deviation of the residuals for ADMM and random search with a RSSM or shooting with a SSM.
-The statistics are taken over the first $30$ iterations of all the ADMM usage over the test period.
-Fig.
-9 shows the mean and standard deviation of the residuals, taken over the first $30$ iterations of each ADMM usage.
-In both cases, the residuals are decreasing and the algorithm shows excellent convergence properties.
-The algorithm converges rapidly in the first few iterations.
-In practice, we found out that performing just 20 iterations of ADMM is a good choice to save computational time.
-It makes little difference when using RSSM because trajectories are sampled beforehand.
-However, with the shooting method, the computation time increases linearly with the number of ADMM iterations.
-It takes one minute to perform ten iterations and ten minutes to perform a hundred iterations on a machine with 8 CPUs (3.5GHz).
-VI Discussion and further work
-The results presented above demonstrate the effectiveness of the algorithm to distribute a global coupled power constraint to local heating systems in a multi-zone building.
-The experiments have revealed that a low prediction error is key to the success of the method.
-Prediction errors are not an issue if one aims at maintaining the power around a target value.
-However, most of the constraints implying a maximum power are hard constraints and an inaccurate forecast will cause the constraint to be violated.
-The maximum prediction error could easily exceed the mean error of $3.72\%$ or $5.81\%$ reached by our models, especially at the end of peak periods where resetting the setpoint changes to zero causes sharp variations of the power usage.
-To improve the accuracy of the models, one may include the state of the neighboring zones in the observations of a zone.
-Another possibility is to retrain the models on data collected during the control period.
-Such data are more representative of encountered states than the data coming from random control of the temperature setpoints used to first train the models.
-An interesting extension of this work would be to measure thermal comfort based on predictions of the temperature of each zone instead of on the temperature setpoints.
-The local objectives $g$, defined in eq.
-(4), would become
-$g(\boldsymbol{\Delta}_{i})=\left\lVert T_{t,i}-T^{sp}_{t,i}\right\rVert_{2}^{2},$
-where $T_{t,i}$ is the temperature in zone $i$ at time $t$.
-This complicates the theoretical analysis because such a $g$ function is no longer convex.
-In addition, some preliminary experiments suggest that it is hard to accurately predict the temperature with fully data-driven models.
-If the temperature setpoint is passed as an input, the models often learn to predict the setpoint as the temperature forecast.
-The use of physics-informed models may be necessary to ensure accurate temperature predictions [23, 37].
-We trained all the prediction models from scratch.
-However, there are many similarities from one zone to the other.
-An interesting improvement for this work will be to leverage the training of a model in one zone to accelerate the training of other models.
-Existing techniques of Meta-Learning could be used for this purpose.
-This will greatly improve the generalization of the approach as models from some buildings could be used for new buildings.
-Finally, tests are all performed using a virtual environment.
-Even though EnergyPlus provides state-of-the-art accuracy for building simulation, future works should include tests on real buildings to further validate the method.
-VII Conclusion
-In this work, we present a distributed algorithm that combines ADMM and Deep Learning models to control the temperature setpoints and act on the power consumption of HVAC systems in buildings.
-Based on a global coupled power constraint, the algorithm distributes the power across the Local Controller systems.
-This approach does not assume any particular type of local heating or cooling system and learns its model entirely from data.
-We provided a theoretical framework and derived a practical algorithm that we implemented to test the efficiency of the approach.
-Tests are performed on a large residential building composed of 18 zones, modeled using EnergyPlus.
-By combining the zones’ prediction models, we reached a mean error of $3.72\%$ with SSM and $5.72\%$ with RSSM on the prediction of the total heating power consumption.
-The Distributed Planning Networks algorithm is tested on a power control task during DR events.
-The power is successfully maintained below a given maximum power limit and the effort is shared across all zones.
-Appendix A Appendix
-A-A Distributed Planning Networks
+(5) ) is important as the building power consumption will be driven towards $\textbf{P}^{tot}$.
+To find the right $\textbf{P}^{tot}$, we leverage the prediction model of each zone.
+We forecast the total power with no setpoint changes ($\Delta=0^{\circ}C)$ to find the business as usual power consumption ($P^{bu}_{t}$) and with the minimum setpoint change ($\delta_{t,i}=m_{i}^{\circ}C$ $\forall t,i$) to find a lower bound on the power consumption ($P^{lb}_{t}$).
+If for all $t$ $P^{bu}_{t}$ is below $P_{t}^{max}$, the optimal solution of the Problem (1) is $\boldsymbol{\Delta}=0$ and no further computation is required.
+If for any $t$, $P^{lb}_{t}$ is above $P_{t}^{max}$, the Problem (1) is unfeasible.
+In the latter case, depending on the energy pricing, one could either reduce as much as possible the temperatures to lower the consumption or take no action to preserve comfort.
+If for all $t$, $P^{max}_{t}$ falls between $P^{bu}_{t}$ and $P^{lb}_{t}$, we proceed with the DPN algorithm and set $\textbf{P}^{tot}=\min(P^{bu}_{t},(1-\nu)P^{max}_{t})$, where $\nu$ is a slack parameter.
+Using a slack is important to satisfy the inequality constraint of Problem (1) with the formulation of Problem (5).
+In addition, as discussed in the next section, it helps mitigate prediction errors.
+This constraint conversion procedure is detailed in Algorithm 3.
 Algorithm 2 Distributed Planning Networks
@@
 27:  end for
+Algorithm 3 Constraint conversion from $\textbf{P}^{max}$ to $\textbf{P}^{tot}$
+1:  Inputs : current state $\text{Inputs : current state }s_{t}\text{ and }\textbf{P}^{max}$
+2:  Prediction models $\text{Prediction models }\phi_{1},\ldots,\phi_{N}$
+3:  Parameters : $\text{Parameters : }\nu,\{m_{i}\}_{i=1,\ldots,N},H$
+4:  Initialize $\textbf{P}^{tot}\in\mathbb{R}^{H}$
+5:  for i=1, …, N do
+6:     $\textbf{P}_{i}^{bu}=\phi_{i}(Q(s_{t},\boldsymbol{0}))$
+7:     $\textbf{P}_{i}^{lb}=\phi_{i}(Q(s_{t},\boldsymbol{\Delta}_{min}))$
+8:  end for
+9:  $\textbf{P}^{bu}=\sum_{i=1}^{N}\textbf{P}_{i}^{bu},\text{ }\textbf{P}^{lb}=\sum_{i=1}^{N}\textbf{P}_{i}^{lb}$
+10:  if $\textbf{P}^{bu}\leq(1-\nu)\textbf{P}^{max}$ then
+11:     $\delta_{t,i}=0\text{ }\forall t,i$
+12:     DO ADMM = FALSE
+13:  else if $\textbf{P}^{lb}>(1-\nu)\textbf{P}^{max}$ then
+14:     $\delta_{t,i}=m_{i}\text{ }\forall t,i$
+15:     DO ADMM = FALSE
+16:  else
+17:     $\textbf{P}^{tot}=\min(\textbf{P}^{bu},(1-\nu)\textbf{P}^{max})$
+18:     DO ADMM = TRUE
+19:  end if
+IV Experiments
+IV-A Description of the test environment
+To test the DPN algorithms, we consider the low-rise apartment building from the set of residential reference buildings developed by the Department of Energy (DOE) [37].
+The building is modeled in EnergyPlus, a software that provides thermal simulations for buildings with state-of-the-art accuracy.
+The building is composed of 3 floors of 6 apartments, each apartment being a thermal zone.
+Apartments are equipped with independent HVAC systems, temperature setpoints schedules, occupancy schedules, and appliance usage schedules.
+In Quebec, the most significant strains on the electricity grid happen during winter, due to cold weather and electric heating.
+We thus consider winter periods ranging from January $1^{st}$ to March $31^{st}$ using Typical Meteorological Year (TMY) files from the city of Montreal.
+The observation, actions, and disturbances are composed of:
+  • observation: zone temperature and HVAC power
+  • disturbances: outdoor temperature, humidity, direct normal irradiance, hour of the day, day of the week.
+  • action: heating temperature setpoint.
+For simplicity, we consider having access to perfect weather prediction.
+This assumption is reasonable because this study only considers short-term predictions on which the weather can be accurately predicted [38, 39].
+IV-B Prediction accuracy of the zones’ models
+(a) 1h ahead predictions
+(b) 2h ahead predictions
+(c) 4h ahead predictions
+Figure 4: Mean Average Error of the SSM and RSSM in each zone for different horizons.
+Each model is trained with 5 different training seeds and a dot represents the mean score on the test set for a training seed.
+Horizon SSM RSSM
+mean min max mean min max
+1h 4.16 3.94 4.30 6.84 6.45 7.19
+2h 6.88 6.54 7.10 7.29 7.00 7.69
+4h 9.70 9.19 10.05 8.92 8.41 9.78
+TABLE I: Mean Average Percentage Error (%) on the entire building HVAC power consumption.
+Predictions are made by aggregating the predictions of each zone.
+The mean, minimum and maximum are taken other the 5 training seeds.
+Figure 5: Example of power predictions compared to the actual power consumption for one zone.
+[…]
+Fig.
+4 shows the Mean Average Error (MAE) of the models for different prediction horizons.
+For 1h ahead predictions, the SMMs perform better than the RSSMs.
+However, as the horizon lengthens the RSSMs give more accurate predictions than the SMMs.
+Note that the magnitude of the score for the models of zones 7 to 11 is lower because these zones are on the middle floor and require less heating.
+As illustrated in Fig.
+5, both SSMs and RSSMs give accurate predictions of power consumption.
+Specifically, spikes in the power consumption corresponding to setpoint changes are captured by both model types.
+Note that these predictions are made without observing the states of the neighboring zones.
+The models’ predictions can be aggregated to forecast the entire building’s HVAC power and Table I shows the Mean Average Percentage Error for both model types.
+As presented in the next section, the aggregated prediction is crucial to verify if the building’s HVAC power satisfies the inequality constraint.
+IV-C DPN control perforamances
+During winter in Quebec, peaks mostly occur in the morning between 6 AM and 9 AM and in the evening between 4 PM and 8 PM [40].
+To test the DPN algorithm we focus on the morning periods and consider DR events every morning of the test period introduced in Section IV-B.
+These events translate in a maximum power constraint of $P^{max}_{t}=21kW$ between $6AM$ and $9AM$.
+To satisfy this constraint, the algorithm must reduce the load by $0\%$ to $25\%$, depending on the day.
+This wide range of power reductions allows us to investigate the ability of the algorithm to handle small adjustments and significant modifications of the zones’ temperature.
+With this setting a total of 10 days out of the 28 days contained events requiring the algorithm to take actions for peak-load reductions.
+For each zone, we set $-2^{\circ}C\leq\delta_{t}\leq 0^{\circ}C$ with increments of $0.25^{\circ}C$.
+The simulation timestep is $15min$ and re-planning occurs every two timesteps to let the zones’ temperature vary after setpoint changes.
+Note that for the Stochastic DPN, the experiment are repeated with $5$ different computation seeds.
+The prediction horizon is set to 1 hour, resulting in 81 choices for setpoint changes.
+With this low-dimension action space, all possible actions are forecasted.
+For each chosen trajectory of setpoint changes, 100 state trajectories are sampled from the RSSM distribution.
+IV-C1 Control performance
+Date Constraint Violation Predicted Constraint Violation
+DDPN SDPN DDPN SDPN
+0% 5% 10% 1 2 2 ($\nu=5\%$) 3 4 5 0% 5% 10% 1 2 2 ($\nu=5\%$) 3 4 5
+02/04 83.3 41.7 0.0 0.0 0.0 0.0 16.7 16.7 16.7 50.0 0.0 0.0 75.0 50.0 0.0 66.7 33.3 83.3
+02/16 33.3 16.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 8.3 0.0 0.0 41.7 33.3 0.0 25.0 33.3 25.0
+02/17 66.7 16.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 41.7 0.0 0.0 41.7 50.0 0.0 41.7 33.3 58.3
+02/18 66.7 25.0 8.3 16.7 16.7 0.0 0.0 0.0 0.0 41.7 0.0 0.0 50.0 58.3 0.0 66.7 58.3 58.3
+02/21 58.3 8.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 58.3 0.0 0.0 33.3 58.3 0.0 58.3 50.0 41.7
+02/22 83.3 25.0 8.3 0.0 0.0 0.0 0.0 8.3 8.3 41.7 0.0 0.0 50.0 83.3 8.3 58.3 66.7 66.7
+02/23 75.0 25.0 8.3 0.0 8.3 0.0 0.0 0.0 8.3 58.3 0.0 0.0 58.3 58.3 8.3 66.7 58.3 58.3
+02/24 33.3 16.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 41.7 0.0 0.0 33.3 33.3 0.0 25.0 33.3 41.7
+02/26 75.0 41.7 8.3 0.0 0.0 0.0 0.0 0.0 0.0 50.0 0.0 0.0 58.3 33.3 0.0 33.3 33.3 33.3
+02/28 25.0 0.0 0.0 0.0 25.0 0.0 8.3 25.0 0.0 0.0 0.0 0.0 8.3 0.0 0.0 8.3 0.0 8.3
+TABLE II: Percentage of timesteps where the constraint is not satisfied, and not satisfied in prediction during planning.
+For DDPN, results are shown for different slack values $\nu$.
+For the SDPN, the slack is set to $\nu=0\%$ and results are displayed for each computation seed.
+(a) Determinisitc DPN ($\nu=10\%)$
+(b) Stochastic DPN
+Figure 6: Three days control example for the Deterministic and Stochastic DPN. For simplicity, the results of a single seed are displayed for the SDPN.
+Figure 7: Effect of the slack parameter $\nu$ of the Deterministic DPN control.
+Figure 8: Mitigation of the rebound peak for the Deterministic DPN algorithm. For the red curve the event ends at 9 AM while for the blue curve it ends at 12 AM.
+Deterministic DPN zone mean $(^{\circ}C)$ $\delta=0$ $0<\delta\leq 1$ $1<\delta\leq 2$ zone mean $(^{\circ}C)$ $\delta=0$ $0<\delta\leq 1$ $1<\delta\leq 2$
+1 -0.72 38 20 42 10 -0.50 39 39 21
+2 -0.47 45 32 23 11 -0.54 33 39 27
+3 -0.72 33 29 38 12 -0.71 41 20 39
+4 -0.86 27 30 42 13 -0.89 35 9 56
+5 -0.58 47 21 32 14 -0.37 39 52 9
+6 -0.44 45 26 29 15 -0.58 45 26 29
+7 -0.48 36 41 23 16 -0.73 23 39 38
+8 -0.52 35 42 23 17 -0.81 35 15 50
+9 -0.81 33 21 45 18 -0.55 36 36 27
+Stochastic DPN zone mean $(^{\circ}C)$ $\delta=0$ $0<\delta\leq 1$ $1<\delta\leq 2$ zone mean $(^{\circ}C)$ $\delta=0$ $0<\delta\leq 1$ $1<\delta\leq 2$
+1 -1.14 12 20 68 10 -1.11 12 17 71
+2 -1.25 14 14 73 11 -0.92 12 33 55
+3 -1.18 11 23 67 12 -1.23 11 12 76
+4 -1.12 15 18 67 13 -1.23 12 20 70
+5 -1.06 15 20 65 14 -1.21 12 15 73
+6 -1.20 12 14 74 15 -1.22 12 18 70
+7 -1.13 15 18 67 16 -1.33 11 09 80
+8 -1.13 14 09 77 17 -1.25 12 15 73
+9 -1.12 15 18 67 18 -1.24 11 12 77
+TABLE III: Setpoint changes distribution for each zone during the DR events requiring actions.
+For the SDPN, results are consistent across all the seeds, for simplicity the results of a single seed are presented.
+For the DDPN, $\nu=10\%$.
+Table II shows the percentage of timesteps where the power constraint was not satisfied for each of the DR events requiring actions.
+When the slack parameter $\nu=0\%$ the algorithm aims at consuming exactly $P^{max}$ when the predicted business-as-usual power exceeds $P^{max}$, which often leads to constraint violations.
+We observe that increasing the slack makes the DDPN algorithm aim below the constraint, which decreases the percentage of constraint violation.
+It is interesting to notice that for the DDPN with slacks of $5\%$ and $10\%$, the constraints are always satisfied in predictions.
+It means that the control algorithm is working perfectly and constraint violations are due to prediction errors only.
+Increasing the slack helps mitigate the prediction errors but also increases the power reduction, thus impacting the comfort.
+A balance must be made between constraint satisfaction and comfort.
+Fig.
+7 illustrates the effect of the slack parameter on the total power consumption.
+For the SDPN performances, the optimization is based on the worst-case scenario and $\nu$ is set to $0\%$ to avoid excessive conservatism.
+The algorithm thus aims at consumption $P^{max}$, which explains the high number of constraint violations in prediction.
+Despite optimizing based on the worst case, the algorithm still aims at consuming exactly the maximum power available, which may lead to constraint violations.
+This can easily mitigated by using non-zero slack.
+To illustrate this, we re-run the second seed that has the most constraint violation with a slack of $5\%$.
+The column ”$2$ $(\nu=5\%)$” in Table II shows that every event was satisfied with this setting.
+Fig.
+6 compares the actual power consumption and the predicted power consumption using the DPN algorithms against the baseline power consumption from the built-in EnergyPlus controller.
+On the first day displayed, the predicted power consumption is around $(1-\nu)P^{max}$ for the DDPN algorithm.
+However, for one timestep, we observe that the consumed power exceeds the maximum.
+This one-timestep prediction error can happen even with excellent average prediction performances and illustrates the importance of prediction accuracy.
+On the same day, we see that the SDPN algorithm takes greater setpoint reductions to satisfy the constraint.
+It is important to notice that our problem formulation with a hard constraint on the power is the most challenging setting because it implies extremely accurate predictions to avoid an over-conservative algorithm.
+That said, depending on the pricing, a hard constraint is not always required.
+Since the DPN algorithm is based on the optimization problem (5) it may also be used as is to effectively limit the power consumption without a hard constraint on the power.
+On the third day displayed in Fig.
+6, we observe the ability of DDPN to slightly adjust the temperature to satisfy the constraint.
+The SDPN algorithm is more conservative and takes greater reductions.
+One possible drawback of using a large slack value is to take unnecessary actions to lower the consumption, as shown on the second day in Fig.
+6(a).
+As illustrated in Fig.
+6, a rebound peak may occur after the DR event because the setpoints are increased at the same time in every zone.
+This can be mitigated by increasing the duration of the power constraint to restart the heating systems later in the day when the total power consumption is lower, as shown in Fig.
+8.
+Regarding the setpoint changes, Table III shows that each zone participates in the power reduction effort.
+While there is nothing explicitly enforcing fairness in our algorithm, the similar mean setpoint changes across the zones may be explained by the uniform initialization of ADMM, which first sets a setpoint change of $-1^{\circ}C$ in each zone.
+We also observe that the setpoint changes are larger with SDPN because of the robust optimization.
+IV-C2 DPN convergence
+Fig.
+9 shows the mean and standard deviation of the residuals, taken over the first 20 iterations of each ADMM usage.
+In both cases, the residuals are decreasing and at each iteration the algorithm converges to a solution.
+Concerning computational time, resolving the LC problems consumes the most time.
+As shown in Table IV, for DDPN, LC computation accounts for nearly 98% of an iteration duration.
+This process is conducted in parallel for each zone and, with sufficient computing resources, is independent of the number of zones, making the algorithm scalable to the size of the building.
+We also notice that the SDPN is significantly faster.
+This is because the action space is low-dimensional and state trajectories are precomputed, as explained in Section III-B2.
+That said, considering longer horizons or larger comfort intervals would increase exponentially the action space dimension.
+Exploring larger action spaces randomly becomes rapidly computationally intractable, unlike the gradient-based search used in DDPN.
+(a) DDPN
+(b) SDPN
+Figure 9: Deterministic and Stochastic DPN convergence
+Deterministic DPN Stochastic DPN
+DPN call 160.0 $\pm$ 42.04 14.43 $\pm$ 0.39
+Coordinator iteration 0.11 $\pm$ 0.02 0.09 $\pm$ 0.01
+LC iteration 3.05 $\pm$ 0.79 0.03 $\pm$ 0.0
+TABLE IV: Mean $\pm$ std of the computation time in seconds for the entire DPN algorithm, the coordinator and local controllers problem.
+Note that each local controller was running on a 2-core CPU with 4GB of RAM.
+V Discussion and further work
+The results presented above demonstrate the effectiveness of the algorithm to distribute a global coupled power constraint to local heating systems in a multi-zone building.
+The experiments have revealed that a low prediction error is key to the success of the method.
+Prediction errors are not an issue if one aims at maintaining the power around a target value.
+However, most of the constraints implying a maximum power are hard constraints and an inaccurate forecast will cause the constraint to be violated.
+The maximum prediction error could easily exceed the mean error of $4.16\%$ or $6.84\%$ reached by our models, especially after setpoint changes when the power consumption varies sharply.
+To improve the accuracy of the models, one may include the state of the neighboring zones in the observations of a zone.
+Another possibility is to retrain the models on data collected during the control period.
+Such data are more representative of encountered states than the data coming from random control of the temperature setpoints used to first train the models.
+An interesting extension of this work would be to measure thermal comfort based on predictions of the temperature of each zone instead of on the temperature setpoints.
+The local objectives $g$, defined in Eq.
+(12), would become
+$g(\boldsymbol{\Delta}_{i})=\left\lVert T_{t,i}-T^{sp}_{t,i}\right\rVert_{2}^{2},$
+where $T_{t,i}$ is the temperature in zone $i$ at time $t$.
+Such a formulation would permit pre-heating strategies and further enhance energy conservation.
+However, it complicates the theoretical analysis because such a $g$ function is no longer convex.
+In addition, some preliminary experiments suggest that it is hard to accurately predict the temperature with fully data-driven models.
+If the temperature setpoint is passed as an input, the models often learn to predict the setpoint as the temperature forecast.
+The use of physics-informed models may be necessary to ensure accurate temperature predictions [23, 41].
+In addition, our experiments showed that relaxing the inequality constraint of Problem (1) with the symmetric term $||A\boldsymbol{u}-\boldsymbol{P}^{tot}||_{2}^{2}$ can ensure the constraint satisfaction given the right choice of $\mathbb{P}^{tot}$ and normalized variables $\boldsymbol{u}$.
+For the algorithm to converge, the function $\ell$ needs to have a Lipschitz continuous gradient so an interesting research direction is to use an asymmetric function such as a log-barrier to relax the inequality constraint.
+The convergence proof will still hold and it could remove the need for a slack parameter.
+However, the coordinator problem will no longer be a simple penalized least-square problem and will be harder to solve.
+We trained all the prediction models from scratch.
+However, there are many similarities from one zone to the other.
+An interesting improvement for this work will be to leverage the training of a model in one zone to accelerate the training of other models.
+Existing techniques of Meta-Learning could be used for this purpose.
+This will greatly improve the generalization of the approach as models from some buildings could be used for new buildings.
+Finally, tests are all performed using a virtual environment.
+Even though EnergyPlus provides state-of-the-art accuracy for building simulation, future works should include tests on real buildings to further validate the method.
+VI Conclusion
+In this work, we present a distributed algorithm that combines ADMM and Deep Learning models to control the temperature setpoints and act on the power consumption of HVAC systems in buildings.
+Based on a global coupled power constraint, the algorithm distributes the power across the Local Controller systems.
+This approach does not assume any particular type of local heating or cooling system and learns its model entirely from data.
+We provided a theoretical framework and derived a practical algorithm that we implemented to test the efficiency of the approach.
+Tests are performed on a large residential building composed of 18 zones, modeled using EnergyPlus.
+By combining the zones’ prediction models, we reached a mean error of $4.16\%$ with SSM and $6.84\%$ with RSSM on the prediction of the total heating power consumption for the $1h$ ahead horizon used for planning.
+Both the deterministic and stochastic versions of the Distributed Planning Networks algorithm are tested on a power control task during DR events.
+The power is successfully reduced below a given maximum power limit and the effort is shared across all zones.
+Appendix A Appendix
+A-A Proof of Theorem 1
+For convenience, let us re-state here the problem.
+Note that to ease the notation we rename the decision variables of the problem by $\boldsymbol{x}$ instead of $\boldsymbol{\Delta}$.
+Recall that for $i=1,\ldots,N$, we have $\boldsymbol{x}_{i}\in\mathbb{R}^{H}$, $\boldsymbol{\bar{x}}_{i}\in\mathbb{R}^{H}$, $g(\boldsymbol{x}_{i})=\|\boldsymbol{x}_{i}\|^{2}_{2}$ and $B_{i}\in\mathbb{R}^{NH\times H}$ such that $\boldsymbol{x}=\displaystyle\sum_{i=1}^{N}B_{i}\boldsymbol{x}_{i}$, where $\boldsymbol{x}\in\mathbb{R}^{NH}$.
+In the following, to ease the notation we drop the subscript to write $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,N}=\{\boldsymbol{x}_{i}\}$.
+Consider the following problem :
+$\underset{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}}{\text{minimize}}\sum_{i=1}^{N}g(\boldsymbol{x}_{i})+\ell\left(\sum_{i=1}^{N}B_{i}\boldsymbol{\bar{x}}_{i}\right)$ (27)
+subject to: (28)
+$\boldsymbol{\bar{x}}_{i}=\boldsymbol{x}_{i},\quad i=1,\ldots,N$ (29)
+$\boldsymbol{x}_{i}\in\chi_{i},\quad i=1,\ldots,N$ (30)
+Where $\chi_{i}$ is some closed convex set. The problem is solved using Algorithm 1. The augmented Lagrangian of this problem is
+($\mathcal{L}(\{\boldsymbol{x}_{i}\},\{\bar{\boldsymbol{x}_{i}}\},\{\boldsymbol{\lambda}_{i}\})=\sum_{i=1}^{N}g(\boldsymbol{x}_{i})+\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{x}}_{i}\right)$ (31)
+$+\sum_{i=1}^{N}\boldsymbol{\lambda}_{i}^{T}(\bar{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i})+\sum_{i=1}^{N}\frac{\rho}{2}\|\bar{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|^{2}_{2}.$
+The proof follows the same steps as in [29], but some differences arise because we are considering $N$ duplicated variables.
+Yet, as we show, the theorem holds mainly because of the orthogonality of the duplicated variables in $\mathbb{R}^{NH}$.
+In addition, note that compared to [29] we consider an update of every variable at each iteration.
+First, we state the following important properties :
+Proposition A.1.
+The function $g$ is convex.
+Proposition A.2.
+For $i=1,\ldots,N$ and $k\in\mathbb{N}$,
+$\|B_{i}\boldsymbol{x}_{i}^{k+1}-B_{i}\boldsymbol{x}_{i}^{k}\|_{2}=\|\boldsymbol{x}_{i}^{k+1}-\boldsymbol{x}_{i}^{k}\|_{2}$.
+Proof.
+The vector $B_{i}\boldsymbol{x}_{i}^{k}$ has zero entries except in the positions $iH$ to $(i+N)H$ where it has the entries of $\boldsymbol{x}_{i}^{k}$. ∎
+Proposition A.3.
+For $k\in\mathbb{N}$,
+$\|\bar{\boldsymbol{x}}^{k+1}-\bar{\boldsymbol{x}}^{k}\|_{2}^{2}=\sum_{i=1}^{N}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|_{2}^{2}$
+Proof.
+The vectors $\{B_{i}\boldsymbol{x}_{i}\}_{i=1,\ldots,N}$ are orthogonal, which yields
+$\|\bar{\boldsymbol{x}}^{k+1}-\bar{\boldsymbol{x}}^{k}\|_{2}^{2}$ $=\left\|\sum_{i=1}^{N}B_{i}(\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k})\right\|_{2}^{2}$ (32)
+$=\sum_{i=1}^{N}\left\|B_{i}(\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k})\right\|_{2}^{2}$ (33)
+$=\sum_{i=1}^{N}\left\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\right\|_{2}^{2}$ (34)
+In addition, we make the following assumptions :
+Assumption A.1.
+The penalty parameter $\rho$ is chosen large enough such that the problems of Algorithm 1 Line 5 and 7 are strongly convex with parameters $\{\gamma_{i}\}$ and $\bar{\gamma}$ respectively.
+See [42] for more details about Assumption A.1.
+Assumption A.2.
+  1. 1.
+    We suppose that $\ell$ has a Lipschitz continuous gradient, i.e.
+there exists a $L>0$ such that for all x and y, $\|\nabla\ell(\textbf{x})-\nabla\ell(\textbf{y})\|_{2}\leq L\|\textbf{x}-\textbf{y}\|_{2}.$
+  2. 2.
+    We suppose $\rho\geq L$ and $\rho\bar{\gamma}\geq 2L^{2}$, where $\bar{\gamma}$ is the strong convexity parameter defined in Assumption A.1.
+Assumption A.3.
+The function $f(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N})=\sum_{i=1}^{N}g(\boldsymbol{x}_{i})+\ell\left(\sum_{i=1}^{N}B_{i}\boldsymbol{x}_{i}\right)$ is lower bounded over $\prod_{i=1}^{N}\chi_{i}$.
+With this framework, we demonstrate the following lemmas:
+Lemma A.1.
+For $i=1,\ldots,N$
+$B_{i}^{T}\nabla\ell\left(\bar{\boldsymbol{x}}\right)=-\boldsymbol{\lambda}_{i}^{k+1}$ (35)
+and
+$L^{2}\|\ \bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|^{2}_{2}\geq\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|^{2}_{2}$ (36)
+where $\nabla_{\bar{\boldsymbol{x}}}\ell\in\mathbb{R}^{NH}$ is the gradient of $\ell$ with respect to $\bar{\boldsymbol{x}}=\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{x}}_{i}$.
+Proof.
+The first order optimality condition of the problem Algorithm 1 Line 7 yields $N$ vector equalities, one for each $\bar{\boldsymbol{x}}_{i}$ variable:
+$\nabla_{\bar{\boldsymbol{x}}_{i}}\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{x}}_{i}\right)+\boldsymbol{\lambda}_{i}^{k}+\rho(\bar{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}^{k+1})=0.$ (37)
+Using the dual update (Algorithm 1 Line 9), the $\boldsymbol{x}_{i}^{k+1}$ are such that
+$\nabla_{\bar{\boldsymbol{x}}_{i}}\ell\left(\sum_{i=1}^{N}B_{i}\bar{\boldsymbol{x}}_{i}^{k+1}\right)=-\boldsymbol{\lambda}_{i}^{k+1}$ (38)
+$B_{i}^{T}\nabla\ell\left(\bar{\boldsymbol{x}}^{k+1}\right)=-\boldsymbol{\lambda}_{i}^{k+1}.$ (39)
+To prove the inequality, we use Assumption A.2 and the equality (35) :
+$\|B_{i}^{T}\nabla\ell\left(\bar{\boldsymbol{x}}^{k+1}\right)-B_{i}^{T}\nabla\ell\left(\bar{\boldsymbol{x}}^{k}\right)\|^{2}_{2}=\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|^{2}_{2}$ (40)
+$L^{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|^{2}_{2}\geq\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|^{2}_{2}$ (41)
+Lemma A.2.
+The difference of two consecutive values of the augmented Lagrangian is bounded by the following negative quantity :
+($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\})$ (42)
+$\leq\sum_{i=1}^{N}\left(\frac{-\gamma_{i}}{2}\|\boldsymbol{x}^{k+1}_{i}-\boldsymbol{x}^{k}_{i}\|^{2}-\left(\frac{\bar{\gamma}}{2}-\frac{L^{2}}{\rho}\right)\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|^{2}\right)$
+Proof.
+The idea is to split the difference of two consecutive values of the augmented Lagrangian to explicit the update of each variable:
+($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\})$ (43)
+=($=\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})$
++($\quad\quad+\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})$
++($\quad\quad+\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\}).$
+First, using the dual update rule (Algorithm 1 Line 9) and Lemma A.1 we have
+($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})$
+$\quad\quad=\frac{1}{\rho}\sum_{i=1}^{N}\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|^{2}_{2}$ (44)
+$\quad\quad\leq\frac{L^{2}}{\rho}\sum_{i=1}^{N}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|^{2}_{2}$ (45)
+In addition, $\mathcal{L}$ is strongly convex (Assumption A.1) so
+($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})$
++($\quad+\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})-\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\})$
+$\leq\sum_{j=1}^{N}\Bigl{(}\bigl{[}\nabla_{\boldsymbol{x}_{j}}\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k}\})\bigr{]}^{T}(\boldsymbol{x}_{j}^{k+1}-\boldsymbol{x}_{j}^{k})$
+$\hskip 113.81102pt-\frac{\gamma_{j}}{2}\|\boldsymbol{x}_{j}^{k+1}-\boldsymbol{x}_{j}^{k}\|^{2}_{2}\Bigr{)}$
+$\hskip 56.9055pt+\Bigl{[}\nabla_{\bar{\boldsymbol{x}}}\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\bar{\boldsymbol{x}},\{\boldsymbol{\lambda}_{i}^{k}\})\Bigr{]}^{T}(\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k})$
+$\hskip 56.9055pt-\frac{\bar{\gamma}_{i}(\rho)}{2}\|\bar{\boldsymbol{x}}^{k+1}-\bar{\boldsymbol{x}}^{k}\|^{2}_{2}$
+$\leq\sum_{i=1}^{N}\Bigl{(}-\frac{\gamma_{i}}{2}\|\boldsymbol{x}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|^{2}_{2}-\frac{\bar{\gamma}}{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|_{2}^{2}\Bigr{)}$ (46)
+The last equality comes from Proposition A.3 and the optimality of the subproblems. Combining the two inequalities (45) and (46) leads to the result. ∎
+Lemma A.3.
+The limit $\lim_{k\rightarrow\infty}\mathcal{L}(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\})$ exists and is bounded from below by $\underset{\boldsymbol{x}_{1},\ldots\boldsymbol{x}_{N}}{min}f(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n})$.
+Proof.
+First, we use the descent property of $\ell$ under the Assumption A.2 :
+$\ell\left(\boldsymbol{x}^{k+1}\right)$ $\leq\ell(\bar{\boldsymbol{x}}^{k+1})+\Bigl{[}\nabla\ell(\bar{\boldsymbol{x}}^{k+1})\Bigr{]}^{T}\Bigl{(}\boldsymbol{x}^{k+1}-\bar{\boldsymbol{x}}^{k+1}\Bigr{)}$
+$\quad\quad+\frac{L}{2}\Bigl{\|}\boldsymbol{x}^{k+1}-\bar{\boldsymbol{x}}^{k+1}\Bigr{\|}^{2}_{2}$
+$\leq\ell(\bar{\boldsymbol{x}}^{k+1})+\sum_{i=1}^{N}\Bigl{(}\Bigl{[}\nabla_{\boldsymbol{\bar{x}}_{i}}\ell(\bar{\boldsymbol{x}}^{k+1})\Bigr{]}^{T}\Bigl{(}\boldsymbol{x}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k+1}\Bigr{)}$
+$\quad\quad+\frac{L}{2}\Bigl{\|}\boldsymbol{x}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k+1}\Bigr{\|}^{2}_{2}\Bigr{)}.$
+Plugging this result in the expression of the Lagrangian yields:
+($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})$
+$\quad=\sum_{i=1}^{N}g(\boldsymbol{x}_{i}^{k+1})+\ell\left(\bar{\boldsymbol{x}}^{k+1}\right)+\sum_{i=1}^{N}\boldsymbol{\lambda}_{i}^{k+1,T}(\bar{\boldsymbol{x}}_{i}^{k+1}-\boldsymbol{x}_{i}^{k+1})$ (47)
+$\hskip 56.9055pt+\sum_{i=1}^{N}\frac{\rho}{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\boldsymbol{x}_{i}^{k+1}\|_{2}^{2}$
+$\quad=\sum_{i=1}^{N}g(\boldsymbol{x}_{i}^{k+1})+\ell\left(\bar{\boldsymbol{x}}^{k+1}\right)$
+$\hskip 56.9055pt+\sum_{i=1}^{N}\left[\nabla_{\boldsymbol{\bar{x}}_{i}}\ell(\bar{\boldsymbol{x}}^{k+1})\right]^{T}\left(\boldsymbol{x}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k+1}\right)$
+$\hskip 56.9055pt+\sum_{i=1}^{N}\frac{\rho}{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\boldsymbol{x}_{i}^{k+1}\|_{2}^{2}$ (48)
+$\quad\geq\sum_{i=1}^{N}g(\boldsymbol{x}_{i}^{k+1})+\ell\left(\boldsymbol{x}^{k+1}\right)+\sum_{i=1}^{N}\frac{\rho-L}{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\boldsymbol{x}_{i}^{k+1}\|_{2}^{2}$ (49)
+$\quad=f(\boldsymbol{x}^{k+1})+\sum_{i=1}^{N}\frac{\rho-L}{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\boldsymbol{x}_{i}^{k+1}\|_{2}^{2}$ (50)
+where (48) comes from Lemma A.1.
+Since $f$ is bounded from below (Assumption A.3), so is ($\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\})$.
+In addition, Lemma A.2 shows that the sequence {($\{\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\}\}_{k}$ is decreasing, which leads to the result.
+∎
+Lemma A.4.
+For $i=1,\ldots,N$,
+$\lim_{k\xrightarrow{}\infty}\|\bar{\boldsymbol{x}}_{i}^{k}-\boldsymbol{x}_{i}^{k}\|_{2}=0$
+Proof.
+We have shown that the sequence {($\{\mathcal{L}(\{\boldsymbol{x}_{i}^{k+1}\},\{\bar{\boldsymbol{x}}_{i}^{k+1}\},\{\boldsymbol{\lambda}_{i}^{k+1}\}\}_{k}$ converges. From Lemma A.2 we have
+$\lim_{k\xrightarrow{}\infty}\|\boldsymbol{x}_{i}^{k+1}-\boldsymbol{x}_{i}^{k}\|_{2}=0$ (51)
+and
+$\lim_{k\xrightarrow{}\infty}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|_{2}=0.$ (52)
+In addition, since $L^{2}\|\bar{\boldsymbol{x}}_{i}^{k+1}-\bar{\boldsymbol{x}}_{i}^{k}\|^{2}_{2}\geq\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|_{2}^{2}$ for all $i$, we have :
+$\lim_{k\xrightarrow{}\infty}\|\boldsymbol{\lambda}_{i}^{k+1}-\boldsymbol{\lambda}_{i}^{k}\|_{2}=0.$ (53)
+The dual update rule (Algorithm 1 Line 9) then ensures that
+$\lim_{k\xrightarrow{}\infty}\|\bar{\boldsymbol{x}}_{i}^{k}-\boldsymbol{x}_{i}^{k}\|_{2}=0.$
+Lemma A.5.
+Any limit point ($(\{\boldsymbol{x}_{i}^{*}\},\{\bar{\boldsymbol{x}}_{i}^{*}\},\{\boldsymbol{\lambda}_{i}^{*}\})$ of the sequence ($(\{\boldsymbol{x}_{i}\},\{\bar{\boldsymbol{x}}_{i}\},\{\boldsymbol{\lambda}_{i}\})$ generated by Algorithm 1 is a stationary solution of the problem (27):
+$(\nabla g(\boldsymbol{x}_{i}^{*})-\boldsymbol{\lambda}_{i}^{*})^{T}(\boldsymbol{y}-\boldsymbol{x}_{i}^{*})\geq 0\quad\forall\boldsymbol{y}_{i}\in\chi_{i}$
+$\nabla_{\boldsymbol{\bar{x}}_{i}}\ell(\bar{\boldsymbol{x}}^{*})-\boldsymbol{\lambda}_{i}^{*}=0,\quad i=1,\ldots,N$
+$\boldsymbol{x}_{i}^{*}=\bar{\boldsymbol{x}}_{i}^{*}\quad i=1,\ldots,N$
+Proof.
+By taking the limit of the result of Lemma A.4, we have that $\boldsymbol{x}_{i}^{*}=\bar{\boldsymbol{x}}_{i}^{*}$ for all $i$.
+In addition, for all $y\in\chi_{i}$ the optimality condition of the problem in Algorithm 1 Line 5 is
+$\left[\nabla_{\boldsymbol{x}_{i}}g(\boldsymbol{x}_{i}^{k+1})-\boldsymbol{\lambda}_{i}^{k}+\rho(\bar{\boldsymbol{x}}_{i}^{k}-\boldsymbol{x}_{i}^{k+1})\right]^{T}(\boldsymbol{y}-\boldsymbol{x}_{i}^{k+1})\geq 0$ (54)
+We previously showed that for all $i$, $\lim_{k\xrightarrow{}\infty}\|\bar{\boldsymbol{x}}_{i}^{k}-\boldsymbol{x}_{i}^{k}\|=0$.
+In addition, Eq.
+(51,52,53) give $\lim_{k\xrightarrow{}\infty}\bar{\boldsymbol{x}}_{i}^{k}=\bar{\boldsymbol{x}}_{i}^{*}$, $\lim_{k\xrightarrow{}\infty}\boldsymbol{x}_{i}^{k}=\boldsymbol{x}_{i}^{*}$, $\lim_{k\xrightarrow{}\infty}\boldsymbol{\lambda}_{i}^{k}=\boldsymbol{\lambda}_{i}^{*}$.
+Taking the limit in the above inequality (54) yields:
+$(\nabla g(\boldsymbol{x}_{i}^{*})-\boldsymbol{\lambda}_{i}^{*})^{T}(\boldsymbol{y}-\boldsymbol{x}_{i}^{*})\geq 0\quad\forall\boldsymbol{y}_{i}\in\chi_{i}.$ (55)
+Taking the limit in the result of Lemma A.2 gives
+$\nabla_{\boldsymbol{\bar{x}}_{i}}\ell(\bar{\boldsymbol{x}}^{*})+\boldsymbol{\lambda}_{i}^{*}=0.$ (56)
+From these lemmas, the proof of Theorem II.1 is identical to the argument given in [29] (Theroem 3.4), and is based on the compactness of the sets $\chi_{i}$.
+For completeness, we remind the theorem and state the argument here.
+Theorem A.1.
+The sequence generated by Algorithm 1 converges to the set of stationary solutions of Problem (14)
+$\lim_{k\xrightarrow{}\infty}dist\left(\{\boldsymbol{x}_{i}^{k}\},\{\bar{\boldsymbol{x}}_{i}^{k}\},\{\boldsymbol{\lambda}_{i}^{k}\};Z^{*}\right)=0,$
+where $Z^{*}$ is the set of primal-dual stationary solutions of the problem. The distance between a vector $\boldsymbol{x}$ and the set $Z^{*}$ is defined as:
+$\text{dist}\left(\boldsymbol{x};Z^{*}\right)=\underset{y\in Z^{*}}{\text{min}}\|x-y\|_{2}$
+Proof.
+First, the compactness of the sets $\chi_{i}$ ensures for all $i$ the sequences $\{\boldsymbol{x}_{i}^{k}\}_{k}$ and $\{\bar{\boldsymbol{x}}_{i}^{k}\}_{k}$ have a limit point.
+In addition, the function $\ell$ has Lipschitz continuous gradients $\nabla_{\bar{\boldsymbol{x}}_{i}}\ell_{i}(\bar{\boldsymbol{x}}_{i}^{k})$, so for all $i$ the set $\{\nabla_{\bar{\boldsymbol{x}}_{i}}\ell_{i}(\bar{\boldsymbol{x}}_{i})\mid\bar{\boldsymbol{x}}_{i}\in\chi_{i}\}$ is bounded and $\{\nabla_{\bar{\boldsymbol{x}}_{i}}\ell_{i}(\bar{\boldsymbol{x}}_{i}^{k})\}_{k}$ is a bounded sequence.
+From Eq.
+(35), $\nabla_{\bar{\boldsymbol{x}}_{i}}\ell_{i}(\bar{\boldsymbol{x}}_{i}^{k})=-\boldsymbol{\lambda}^{k}_{i}$, so the sequences $\{\boldsymbol{\lambda}_{i}^{k}\}_{k}$ also have a limit point.
+The sequences $\{\boldsymbol{x}_{i}^{k}\}_{k}$, $\{\bar{\boldsymbol{x}}_{i}^{k}\}_{k}$, and $\{\boldsymbol{\lambda}_{i}^{k}\}_{k}$ lie in compact sets, so for all $i$ every subsequence ($(\{\boldsymbol{x}_{i}^{k_{j}}\},\{\bar{\boldsymbol{x}}_{i}^{k_{j}}\},\{\boldsymbol{\lambda}_{i}^{k_{j}}\})$ has a limit point ($(\{\hat{\boldsymbol{x}}_{i}\},\{\hat{\bar{\boldsymbol{x}}}_{i}\},\{\hat{\boldsymbol{\lambda}}_{i}\})$.
+From Lemma A.5 we know this limit point is in $Z^{*}$.
+By further restricting the subsequence if necessary, we can assume this limit point is unique.
+Suppose the sequence ($(\{\boldsymbol{x}_{i}^{k_{j}}\},\{\bar{\boldsymbol{x}}_{i}^{k_{j}}\},\{\boldsymbol{\lambda}_{i}^{k_{j}}\})$ does not converge to $Z^{*}$. Then
+$\lim_{j\rightarrow\infty}\text{dist}\left((\{\boldsymbol{x}_{i}^{k_{j}}\},\{\bar{\boldsymbol{x}}_{i}^{k_{j}}\},\{\boldsymbol{\lambda}_{i}^{k_{j}}\});Z^{*}\right)=e>0.$ (57)
+But there exists an $n>0$ such that for all $j\geq n$,
+($\left\|(\{\boldsymbol{x}_{i}^{k_{j}}\},\{\bar{\boldsymbol{x}}_{i}^{k_{j}}\},\{\boldsymbol{\lambda}_{i}^{k_{j}}\})-(\{\hat{\boldsymbol{x}}_{i}\},\{\hat{\bar{\boldsymbol{x}}}_{i}\},\{\hat{\boldsymbol{\lambda}}_{i}\})\right\|_{2}^{2}\leq\frac{e}{2},$ (58)
+and by definition of the distance,
+$\text{dist}\left((\{\boldsymbol{x}_{i}^{k_{j}}\},\{\bar{\boldsymbol{x}}_{i}^{k_{j}}\},\{\boldsymbol{\lambda}_{i}^{k_{j}}\});Z^{*}\right)\leq\frac{e}{2},$ (59)
+which contradicts Eq.
+(57).
+We have shown that every sequence of iterates has a limit point that is a stationary solution.
+Furthermore, the sequence of iterates generated by the algorithm converges to one of these limit points.
+The sequence of iterates produced by Algorithm 1 therefore converges to the set of stationary solutions.
+©2024 IEEE.
+Personal use of this material is permitted.
+Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.